### v3

# Algorithmic contiguity from low-degree conjecture and applications in correlated random graphs

Abstract
In this paper, assuming a natural strengthening of the low-degree conjecture, we provide evidence of computational hardness for two problems: (1) the (partial) matching recovery problem in the sparse correlated Erdős-Rényi graphs $\mathcal{G}(n,q;\rho)$ when the edge-density $q=n^{-1+o(1)}$ and the correlation $\rho<\sqrt{\alpha}$ lies below the Otter’s threshold, this resolves a remaining problem in [DDL23+]; (2) the detection problem between a pair of correlated sparse stochastic block model $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ and a pair of independent stochastic block models $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$ when $\epsilon^{2}\lambda s<1$ lies below the Kesten-Stigum (KS) threshold and $s<\sqrt{\alpha}$ lies below the Otter’s threshold, this resolves a remaining problem in [CDGL24+].
One of the main ingredient in our proof is to derive certain forms of algorithmic contiguity between two probability measures based on bounds on their low-degree advantage. To be more precise, consider the high-dimensional hypothesis testing problem between two probability measures $\mathbb{P}$ and $\mathbb{Q}$ based on the sample $\mathsf{Y}$. We show that if the low-degree advantage $\mathsf{Adv}_{\leq D}\big{(}\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=O(1)$, then (assuming the low-degree conjecture) there is no efficient algorithm $\mathcal{A}$ such that $\mathbb{Q}(\mathcal{A}(\mathsf{Y})=0)=1-o(1)$ and $\mathbb{P}(\mathcal{A}(\mathsf{Y})=1)=\Omega(1)$. This framework provides a useful tool for performing reductions between different inference tasks.
1 Introduction
Graph matching, also referred to as network alignment, is the problem of identifying a bijection between the vertex sets of two graphs that maximizes the number of common edges. When the two graphs are exactly isomorphic to each other, this problem reduces to the classical graph isomorphism problem, for which the best known algorithm runs in quasi-polynomial time [Bab16]. In general, graph matching is an instance of the quadratic assignment problem [BCPP98], which is known to be NP-hard to solve or even approximate [MMS10]. Motivated by real-world applications (such as social network de-anonymization [NS08] and computational biology [SXB08]) as well as the need to understand the average-case computational complexity, recent research has focused on developing theoretical foundations and efficient algorithms for graph matching under statistical models. These models assume that the two graphs are randomly generated with correlated edges under a hidden vertex correspondence, and a canonical model among them is the following correlated random graph model. For any integer $n$, denote by $\operatorname{U}=\operatorname{U}_{n}$ the set of unordered pairs $(i,j)$ with $1\leq i\neq j\leq n$.
Definition 1.1 (Correlated random graph model).
Given an integer $n\geq 1$, for $(i,j)\in\operatorname{U}$ let $J_{i,j}$ and $K_{i,j}$ be independent Bernoulli variables with parameter $s$. In addition, let $\pi_{*}$ be an independent uniform permutation on $[n]=\{1,\dots,n\}$. Then, we define a triple of correlated random graphs $(G,A,B)$ as follows: first we generate $G$ independently with $\{I_{i,j},K_{i,j}\}$ and $\pi_{*}$ from a specific probability distribution over all graphs on $[n]$, and then (conditioned on $G$) we define for each $(i,j)\in\operatorname{U}$ that (note that we identify a graph with its adjacency matrix)
$A_{i,j}=G_{i,j}J_{i,j}\,,\quad B_{i,j}=G_{i,j}K_{\pi_{*}^{-1}(i),\pi_{*}^{-1}(j)}\,.$
In short, we will subsample $A,B$ from $G$ with subsampling probability $s$ and then permute the vertices of $B$ by a uniform permutation.
Of particular interest in our paper are the following two special cases, namely the correlated Erdős-Rényi model and the correlated stochastic block model (SBM).
Definition 1.2 (Correlated Erdős-Rényi graph model).
Given an integer $n\geq 1$ and two parameters $p,s\in(0,1)$, we generate a triple of correlated random graphs $(G,A,B)$ such that we first generate $G$ according to an Erdős-Rényi graph distribution $\mathcal{G}(n,p)$ (i.e., for each $(i,j)\in\operatorname{U}$ we connect $(i,j)$ in $G$ independently with probability $p$), and then generate $(A,B)$ from $G$ according to Definition 1.1. For ease of presentation, we shall reparameterize such that $q=ps$ and $\rho=\frac{s(1-p)}{1-ps}$ respectively. We will denote the marginal law of $(A,B)$ as $\mathcal{G}(n,q;\rho)$.
Definition 1.3 (Stochastic block model).
Given an integer $n\geq 1$ and three parameters $k\in\mathbb{N},\lambda>0,\epsilon\in(0,1)$, we define a random graph $G$ as follows: (1) sample a labeling $\sigma_{*}\in[k]^{n}=\{1,\ldots,k\}^{n}$ uniformly at random; (2) for every distinct pair $(i,j)\in\operatorname{U}_{n}$, we let $G_{i,j}$ be an independent Bernoulli variable such that $G_{i,j}=1$ (which represents that there is an undirected edge between $i$ and $j$) with probability $\frac{(1+(k-1)\epsilon)\lambda}{n}$ if $\sigma_{*}(i)=\sigma_{*}(j)$ and with probability $\frac{(1-\epsilon)\lambda}{n}$ if $\sigma_{*}(i)\neq\sigma_{*}(j)$. In this case, we say that $G$ is sampled from a stochastic block model $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon)$.
Definition 1.4 (Correlated stochastic block model).
Given an integer $n\geq 1$ and four parameters $k\in\mathbb{N},\lambda>0,\epsilon,s\in(0,1)$, we define a triple of correlated random graphs $(G,A,B)$ as follows: we first sample $G$ according to the law of a stochastic block model $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon)$ and then generate $(A,B)$ from $G$ according to Definition 1.1. We will denote the marginal law of $(A,B)$ as $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$.
Two fundamental problems in the study of correlated random graph model are as follows: (1) the detection problem, which involves determining whether a given pair of graphs $(A,B)$ is sampled from a pair of correlated random graphs or from a pair of independent random graphs; (2) the matching problem, which focuses on recovering the latent matching $\pi_{*}$ from a sample $(A,B)$ from the distribution of correlated random graphs. In recent years, significant progress has been made in understanding these problems for both the correlated Erdős-Rényi model and correlated stochastic block models. Through the collective efforts of the community, the information-theoretic thresholds for detection and matching have been fully characterized for correlated Erdős-Rényi model [CK17+, HM23, WXY22, WXY23, GML21, DD23a, DD23b, Du25+] and partially characterized for correlated SBMs [RS21, GRS22]. Additionally, various efficient detecting and matching algorithms have been developed with performance guarantees [DCKG19, BCL+19, DMWX21, FMWX23a, FMWX23b, GM20, GML24, GMS24, MRT21, MRT23, MWXY24, MWXY23, DL22+, DL23+, CR24, CDGL24+, Li24+]. Notably, like many other inference tasks in high-dimensional statistics [ZK16, RSS19, KWB22, Gam21], these problems appear to exhibit information-computation gaps. Specifically, for certain ranges of the correlation strength, detection or matching is information theoretically possible but no efficient algorithm is known to achieve these tasks. We now focus on the algorithmic side of these problems as they are more relevant to our work. Indeed, it has been shown that many inference tasks in the correlated random graph models exhibits sharp algorithmic phase transitions, as we elaborate below:
  • For the detection problem between a pair of correlated Erdős-Rényi models $\mathcal{G}(n,q;\rho)$ and two independent Erdős-Rényi models $\mathcal{G}(n,q)$, we focus on the sparse regime where $q=n^{-1+o(1)}$. In this regime, on the one hand, it was shown in [MWXY24] that when $\rho>\sqrt{\alpha}$ where $\alpha\approx 0.338$ is the Otter’s constant [Ott48], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was shown in [DDL23+] that when $\rho<\sqrt{\alpha}$ there are evidence suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
  • For the detection problem between a pair of correlated SBMs $\mathcal{S}(n,\frac{\lambda}{n};k,\epsilon;s)$ and two independent Erdős-Rényi graphs $\mathcal{G}(n,\frac{\lambda s}{n})$, we focus on the constant degree regime where $\lambda=O(1)$. In this regime, on the one hand, it was shown in [CDGL24+] that when $s>\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ where $\alpha\approx 0.338$ is the Otter’s constant and $\tfrac{1}{\epsilon^{2}\lambda}$ is the Kesten-Stigum threshold [KS66], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was also shown in [CDGL24+] that when $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ there are evidence suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
The lower bound in the aforementioned results explored inherent computational barriers from the perspective of the low-degree polynomial framework. Indeed, it has been proved that the class of low-degree polynomial algorithms is a useful proxy for computationally efficient algorithms, in the sense that the best-known polynomial-time algorithms for a wide variety of high-dimensional inference problems are captured by the low-degree class; see e.g. [HKP+17, Hop18, SW22, KWB22]. However, these aforementioned results suffer from two significant limitations, which we now discuss. Firstly, the aforementioned result cannot provide the evidence that partial matching recovery (that is, recover a positive fraction of the coordinates of the latent matching $\pi_{*}$) is impossible by efficient algorithms although efficient detection has already been ruled out. Secondly, (in the case of correlated SBM) they are only able to establish the computation threshold on the detection problem between the correlated model and a pair of independent Erdős-Rényi graphs. These limitations motivate two natural questions:
Question 1.5.
  1. (1)
    Can we provide the evidence that partial matching recovery is impossible (by efficient algorithms) in the same parameter regime where there are evidence suggesting that detection is impossible (by efficient algorithms)?
  2. (2)
    What can we say about the (arguably more natural) detection problem between a pair of correlated SBMs $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ and a pair of independent SBMs $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$?
The aim of this paper is to (partially) answer these two problems. Our main result can be informally summarized as follows:
Theorem 1.6 (Informal).
We have the following results hold.
  1. (1)
    Assuming a natural strengthening of the low-degree conjecture (see Conjecture 2.5 for its precise meaning), then for the correlated Erdős-Rényi model $\mathcal{G}(n,q,\rho)$, when $q=n^{-1+o(1)}$ and $\rho<\sqrt{\alpha}$ it is impossible to recover a positive fraction of the coordinates of $\pi_{*}$ with constant probability by efficient algorithms.
  2. (2)
    Assuming a natural strengthening of the low-degree conjecture (see Conjecture 2.5 for its precise meaning), then for the correlated stochastic block models $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$, when $\lambda=O(1)$ and $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ it is impossible to recover a positive fraction of the coordinates of $\pi_{*}$ with probability tending to $1$ as $n\to\infty$ by efficient algorithms.
  3. (3)
    Assuming the low-degree conjecture (see Conjecture 2.2 for its precise meaning), then for the correlated stochastic block models $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$, when $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ it is impossible to strongly distinguish this model and a pair of independent SBMs $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$ by efficient algorithms, provided that the average degree $\lambda s$ is sufficiently large.
Remark 1.7.
Note that when $\rho>\sqrt{\alpha}$, the results in [MWXY23, GM20, GML24, GMS24] show that there exists an efficient algorithm that achieves partial recovery of $\pi_{*}$ in a pair of correlated Erdős-Rényi models with probability tending to $1$ as $n\to\infty$. Thus, Item (1) in Theorem 1.6 is tight in some sense and the algorithmic partial recovery threshold is indeed given by $\rho=\sqrt{\alpha}$.
Similarly, the results in [MWXY23, GM20, GML24, GMS24] naturally extends to show that there exists an efficient algorithm that achieves partial recovery of $\pi_{*}$ in a pair of correlated SBMs when $s>\sqrt{\alpha}$ with probability tending to $1$ as $n\to\infty$. Thus, Item (2) in Theorem 1.6 is also tight in the subcritical regime (i.e., when marginally two graphs are below the KS threshold).
Remark 1.8.
From Item (3) in Theorem 1.6 we see that for a pair of correlated SBMs $(A,B)$, when marginally both $A$ and $B$ are below the KS threshold (i.e., when $\epsilon^{2}\lambda s<1$), there is evidence suggesting that no efficient algorithm can strongly distinguish $(A,B)$ from a pair of independent stochastic block models when the correlation $s<\sqrt{\alpha}$. Since the result in [MWXY24] extends naturally to the case of stochastic block models which provides an efficient algorithm that strongly distinguish these two models when $s>\sqrt{\alpha}$, we see that in the subcritical regime (i.e., when marginally both $A$ and $B$ are below the KS threshold) the algorithmic correlation detection threshold is given by $\sqrt{\alpha}$. On the contrary, in the supercritical regime where $A$ and $B$ are above the KS threshold (i.e, when $\epsilon^{2}\lambda s>1$), we believe that the algorithmic correlation detection threshold should be strict lower than $\sqrt{\alpha}$.
1.1 Key challenges and innovations
In this subsection, we briefly discuss our approach of proving Theorem 1.6 and some conceptual innovations behind it. Our idea can be informally summarized as follows:
  1. (1)
    As for Items (1) and (2), for simplicity we take correlated Erdős-Rényi model for example. Denote $\mathbb{P}$ to be the law of $\mathcal{G}(n,q,\rho)$ and $\mathbb{Q}$ to be the law of two independent $\mathcal{G}(n,q)$. We will argue by contradiction and assume that there is a partial recovery algorithm. Then we show that we can use this algorithm to efficiently construct a family of statistics $\{g_{i}:1\leq i\leq n\}$ such that $g_{i}$ approximates $\mathbf{1}_{\{\pi_{*}(1)=i\}}$ under $\mathbb{P}$ in a certain sense.
  2. (2)
    We will show that the low-degree advantage between $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ is bounded by an absolute constant. Then, from the standard low-degree conjecture (see Conjecture 2.2 for details) these two measures cannot be strongly distinguished by efficient algorithms. Thus, since $g_{i}$ is “of positive constant order” under the measure $\mathbb{P}(\cdot\mid\pi_{*}(i)=1)$ (as it should approximate $\mathbf{1}_{\{\pi_{*}(1)=i\}}$ in some sense) we expect that $g_{i}$ should also be “of positive constant order” under $\mathbb{Q}$.
  3. (3)
    Define $g=g_{1}+\ldots+g_{n}$. We then have a statistics $g$ that can be efficiently computed, and (I) under $\mathbb{Q}$ we expect that $g$ is “large” as it is the sum of $n$ “of positive constant order” terms; (II) under $\mathbb{P}$ we expect that $g$ is “not large” as each $g_{i}$ should approximate $\mathbf{1}_{\{\pi_{*}(1)=i\}}$ in some sense. Thus, the statistics $g$ accumulates more signals than all low-degree polynomials, which (non-rigorously speaking) violates the low-degree conjecture.
  4. (4)
    As for Item (3), denote $\mathbb{P}$ to be the law of correlated SBMs and $\mathbb{Q}$ of independent SBMs. Also denote $\widetilde{\mathbb{Q}}$ to be the law of independent Erdős-Rényi graphs. As it was already shown in [CDGL24+] that $\mathbb{P}$ and $\widetilde{\mathbb{Q}}$ cannot be strongly distinguished, (non-rigorously speaking) we only need to show that $\mathbb{Q}$ and $\widetilde{\mathbb{Q}}$ are also “indistinguishable” in some sense.
However, there are certain obstacles when implementing the above ideas, as we shall discuss below. In Step (2) we need to “transfer” the behavior of a statistic $g_{i}$ under $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ to its behavior under $\mathbb{Q}$. If the low-degree advantage between $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ is bounded by $1+o(1)$, then the standard low-degree conjecture implies that we cannot distinguish between $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ better than random by efficient algorithms. Consequently, we expect the behavior of $g_{i}$ to be “almost identical” under both $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ (as $g_{i}$ can be efficiently computed). However, in our case the low-degree advantage is just bounded by a large (but fixed) constant. This weaker condition means that the standard low-degree conjecture only rules out the possibility to distinguish these two measures efficiently with vanishing errors, leaving room for non-negligible distinctions. Similar issues also arise in Step (4), in which simply showing the low-degree advantage between $\mathbb{P}$ and $\widetilde{\mathbb{Q}}$ and the low-degree advantage between $\mathbb{Q}$ and $\widetilde{\mathbb{Q}}$ is bounded is not strong enough for our goal.
To address these issues, one of the main conceptual contributions in our work is to give a more refined characterization of the limitations of efficient algorithms when the low-degree advantage is $O(1)$. Specifically, we show that (assuming low-degree conjecture) bounded low-degree advantage does not only exclude all algorithms that strongly distinguish $\mathbb{P}$ and $\mathbb{Q}$, but also suggest a certain kind of algorithmic contiguity. To be more precise, if the low-degree advantage $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}$ is bounded, then there is no efficient algorithm $\mathcal{A}$ such that $\mathbb{Q}(\mathcal{A}=0)=1-o(1)$ and $\mathbb{P}(\mathcal{A}=1)=\Omega(1)$. This framework allows us to transfer the behavior of a efficiently computable statistics $g$ under different probability measures more easily. For example, if we know that $\mathbb{P}(g\geq c)\geq\Omega(1)$ it immediately holds that $\mathbb{Q}(g\geq c)\geq\Omega(1)$.
Another difficulty is that reducing between different inference tasks requires a framework that better constrains the behavior of efficiently computable estimators. To address this, we introduce a natural refinement of the low-degree conjecture: we posit that low-degree polynomials achieve signal-to-noise ratios that are at least as good as those of any efficient algorithm (see Conjecture 2.5 for a more precise description). While similar modifications to the low-degree conjecture have been explored in [MW23+], our motivations differ significantly. Their work focused on characterizing the limits of precise error rate for all computationally feasible algorithms, whereas our goal is to enable reductions between distinct inference tasks.
1.2 Comparison to concurrent work
In a concurrent work [DHSS25+], the authors investigate the computational hardness of the weak recovery problem in the stochastic block model $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon)$ when $k=n^{o(1)}$ and $\epsilon^{2}\lambda<1$ lies below the KS threshold. Notably, their proof is also based on a “recovery-to-detection reduction” approach, which argues that if there exists an efficient algorithm that achieves weak recovery, then we can use this algorithm to efficiently construct a statistics for detection (in a certain sense). Their results also relies on a strengthening of the low-degree conjecture, which is similar to the assumptions made in this paper.
However, the reduction technique employed in this paper differs significantly from that of [DHSS25+]. The authors of [DHSS25+] construct their statistic for detection using a correlation preserving projection technique established in [HS17]. This technique, which is designed specifically for weak recovery in block models, enables them to “regularize” the statistics and thus bound its moments under the null hypothesis. In contrast, our construction is more straightforward as we leverage the universal framework described in Section 1.1 to “transfer” the behavior of statistics across different probability measures. It seems that in comparison to [DHSS25+], our approach is easier to implement, less dependent on problem-specific details, and potentially applicable to a broader range of problems.
1.3 Notations
In this subsection, we record a list of notations that we shall use throughout the paper. Denote $\mathfrak{S}_{n}$ the set of permutations over $[n]$ and denote $\mu$ the uniform distribution on $\mathfrak{S}_{n}$. In addition, denote $\nu$ to be the uniform distribution on $[k]^{n}$. We will use the following notation conventions on graphs.
  • Labeled graphs. Denote by $\mathcal{K}_{n}$ the complete graph with vertex set $[n]$ and edge set $\operatorname{U}_{n}$. For any graph $H$, let $V(H)$ denote the vertex set of $H$ and let $E(H)$ denote the edge set of $H$. We say $H$ is a subgraph of $G$, denoted by $H\subset G$, if $V(H)\subset V(G)$ and $E(H)\subset E(G)$. Define the excess of the graph $\tau(H)=|E(H)|-|V(H)|$.
  • Isolated vertices. For $u\in V(H)$, we say $u$ is an isolated vertex of $H$, if there is no edge in $E(H)$ incident to $u$. Denote $\mathcal{I}(H)$ the set of isolated vertices of $H$. For two graphs $H,S$, we denote $H\ltimes S$ if $H\subset S$ and $\mathcal{I}(S)\subset\mathcal{I}(H)$, and we denote $H\Subset S$ if $H\subset S$ and $\mathcal{I}(H)=\emptyset$. For any graph $H\subset\mathcal{K}_{n}$, let $\widetilde{H}$ be the subgraph of $H$ induced by $V(H)\setminus\mathcal{I}(H)$.
  • Graph intersections and unions. For $H,S\subset\mathcal{K}_{n}$, denote by $H\cap S$ the graph with vertex set given by $V(H)\cap V(S)$ and edge set given by $E(H)\cap E(S)$. Denote by $S\cup H$ the graph with vertex set given by $V(H)\cup V(S)$ and edge set $E(H)\cup E(S)$. In addition, denote by $S\Cap H$, $S\Cup H$ and S
    H
[…]
    the graph induced by the edge set $E(S)\cap E(H)$, $E(S)\cup E(H)$ and $E(S)\triangle E(H)$, respectively (in particular, these induced graphs have no isolated points).
  • Paths. We say a subgraph $H\subset\mathcal{K}_{n}$ is a path with endpoints $u,v$ (possibly with $u=v$), if there exist distinct $w_{1},\ldots,w_{m}\neq u,v$ such that $V(H)=\{u,v,w_{1},\ldots,w_{m}\}$ and $E(H)=\{(u,w_{1}),(w_{1},w_{2})\ldots,(w_{m},v)\}$. We say $H$ is a simple path if its endpoints $u\neq v$. Denote $\operatorname{EndP}(P)$ as the set of endpoints of a path $P$.
  • Cycles and independent cycles. We say a subgraph $H$ is an $m$-cycle if $V(H)=\{v_{1},\ldots,v_{m}\}$ and $E(H)=\{(v_{1},v_{2}),\ldots,(v_{m-1},v_{m}),(v_{m},v_{1})\}$. For a subgraph $K\subset H$, we say $K$ is an independent $m$-cycle of $H$, if $K$ is an $m$-cycle and no edge in $E(H)\setminus E(K)$ is incident to $V(K)$. Denote by $\mathtt{C}_{m}(H)$ the set of $m$-cycles of $H$ and denote by $\mathcal{C}_{m}(H)$ the set of independent $m$-cycles of $H$. For $H\subset S$, we define $\mathfrak{C}_{m}(S,H)$ to be the set of independent $m$-cycles in $S$ whose vertex set is disjoint from $V(H)$. Define $\mathfrak{C}(S,H)=\cup_{m\geq 3}\mathfrak{C}_{m}(S,H)$.
  • Leaves. A vertex $u\in V(H)$ is called a leaf of $H$, if the degree of $u$ in $H$ is $1$; denote $\mathcal{L}(H)$ as the set of leaves of $H$.
  • Graph isomorphisms and unlabeled graphs. Two graphs $H$ and $H^{\prime}$ are isomorphic, denoted by $H\cong H^{\prime}$, if there exists a bijection $\pi:V(H)\to V(H^{\prime})$ such that $(\pi(u),\pi(v))\in E(H^{\prime})$ if and only if $(u,v)\in E(H)$. Denote by $\mathcal{H}$ the isomorphism class of graphs; it is customary to refer to these isomorphic classes as unlabeled graphs. Let $\operatorname{Aut}(H)$ be the number of automorphisms of $H$ (graph isomorphisms to itself).
For two real numbers $a$ and $b$, we let $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$. We use standard asymptotic notations: for two sequences $a_{n}$ and $b_{n}$ of positive numbers, we write $a_{n}=O(b_{n})$, if $a_{n}<Cb_{n}$ for an absolute constant $C$ and for all $n$ (similarly we use the notation $O_{h}$ is the constant $C$ is not absolute but depends only on $h$); we write $a_{n}=\Omega(b_{n})$, if $b_{n}=O(a_{n})$; we write $a_{n}=\Theta(b_{n})$, if $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$; we write $a_{n}=o(b_{n})$ or $b_{n}=\omega(a_{n})$, if $a_{n}/b_{n}\to 0$ as $n\to\infty$. In addition, we write $a_{n}\circeq b_{n}$ if $a_{n}=[1+o(1)]b_{n}$. For a set $\mathsf{A}$, we will use both #$\#\mathsf{A}$ and $|\mathsf{A}|$ to denote its cardinality. For two probability measures $\mathbb{P}$ and $\mathbb{Q}$, we denote the total variation distance between them by $\operatorname{TV}(\mathbb{P},\mathbb{Q})$.
1.4 Organization of this paper
The rest part of this paper is organized as follows: in Section 2 we state the precise framework of low-degree polynomials and how we relate it to the notion of algorithmic contiguity (see Theorem 2.4). In Section 3 we use this framework to deduce the hardness of partial matching in correlated Erdős-Rényi models and correlated SBMs, thus verifying Items (1) and (2) in Theorem 1.6 (see Theorem 3.2). In Section 4 we again use this framework to deduce the hardness of testing correlated SBMs against independent SBMs, thus verifying Item (3) in Theorem 1.6 (see Corollary 4.2). Several auxiliary results are moved to the appendix to ensure a smooth flow of presentation.
2 Low-degree framework and algorithmic contiguity
The low-degree polynomial framework first emerged from the works of [BHK+19, HS17, HKP+17, Hop18] and has since been refined and extended in various directions. It has found applications across a broad spectrum of problems, including detection tasks such as planted clique, planted dense subgraph, community detection, sparse PCA, and tensor PCA (see [HS17, HKP+17, Hop18, KWB22, SW22, DMW23+, BKW20, DKW+22, MW25, DDL23+, KMW24]), optimization problems like finding maximal independent sets in sparse random graphs [GJW20, Wein22], and constraint satisfaction problems such as random $k$-SAT [BH22]; see also the survey [KWB22]. Additionally, it is conjectured in [Hop18] that the failure of degree-$D$ polynomials implies the failure of all “robust” algorithms with running time $n^{\widetilde{O}(D)}$ (here $\widetilde{O}$ means having at most this order up to a $\operatorname{poly}\log n$ factor). In the remaining of this paper, we will focus on applying this framework in the context of high-dimensional hypothesis testing problems.
To be more precise, consider the hypothesis testing problem between two probability measures $\mathbb{P}$ and $\mathbb{Q}$ based on the sample $\mathsf{Y}\in\mathbb{R}^{N}$. We will be especially interested in asymptotic settings where $N=N_{n},\mathbb{Q}=\mathbb{Q}_{n},\mathbb{P}=\mathbb{P}_{n},\mathsf{Y}=\mathsf{Y}_{n}$ scale with $n$ as $n\to\infty$ in some prescribed way. The standard low-degree polynomial framework primarily focus on the following notions on strong and weak detection.
Definition 2.1 (Strong/weak detection).
We say an algorithm $\mathcal{A}$ that takes $\mathsf{Y}$ as input and outputs either 00 or $1$ achieves
  • strong detection, if the sum of type-I and type-II errors $\mathbb{Q}(\mathcal{A}(\mathsf{Y})=1)+\mathbb{P}(\mathcal{A}(\mathsf{Y})=0)$ tends to 00 as $n\to\infty$.
  • weak detection, if the sum of type-I and type-II errors is uniformly bounded above by $1-\epsilon$ for some fixed $\epsilon>0$.
Roughly speaking, the low-degree polynomial approach focuses on understanding the capabilities and limitations of algorithms that can be expressed as low-degree polynomial functions of the input variables (in our case, the entries of $\mathsf{Y}$). To be more precise, let $\mathcal{P}_{D}=\mathcal{P}_{n,D}$ denote the set of polynomials from $\mathbb{R}^{N}$ to $\mathbb{R}$ with degree no more than $D$. With a slight abuse of notation, we will often refer to “a polynomial” to mean a sequence of polynomials $f=f_{n}\in\mathcal{P}_{n,D}$, where each $f_{n}$ corresponds to problem size $n$; the degree $D=D_{n}$ of such a polynomial may scale with $n$. As suggested by [Hop18], the key quantity is the low-degree advantage
${}\mathsf{Adv}_{\leq D}\Big{(}\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\Big{)}:=\sup_{f\in\mathcal{P}_{D}}\frac{\mathbb{E}_{\mathbb{P}}[f]}{\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}}\,.$ (2.1)
Note that if we denote the likelihood ratio $L(\mathsf{Y})=\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}(\mathsf{Y})$, it is well-known (see, e.g., [Hop18]) the right hand side equals the $L_{2}$-norm of the projection of $L(\mathsf{Y})$ into the subspace spanned by all polynomials of degree bounded by $D$ (the norm is induced by natural inner product under $\mathbb{Q}$) and thus we characterize it with $\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}$. The low degree conjecture, proposed in [Hop18], can be summarized as follows.
Conjecture 2.2 (Low-degree conjecture).
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, the following statements hold:
  1. (1)
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=O(1)$ as $n\to\infty$, then there exists a constant $C$ such that there is no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves strong detection between $\mathbb{P}$ and $\mathbb{Q}$.
  2. (2)
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=1+o(1)$ as $n\to\infty$, then there exists a constant $C$ such that there is no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves weak detection between $\mathbb{P}$ and $\mathbb{Q}$.
Motivated by [Hop18, Hypothesis 2.1.5 and Conjecture 2.2.4] as well as the fact that low-degree polynomials capture the best known algorithms for a wide variety of statistical inference tasks, Conjecture 2.2 appears to hold for distributions of a specific form that frequently arises in high-dimensional statistics. For further discussion on what types of distributions are suitable for this framework, we refer readers to [Hop18, KWB22, Kun21, ZSWB22]. In addition, we point out that although in most applications (and in the statement of [Hop18, Hypothesis 2.1.5]) it is typically to take $\mathbb{Q}$ to be a “null” measure and $\mathbb{P}$ to be a “planted” measure (which makes (2.1) more tractable), several recent works [RSWY23, KVWX23] showed that this framework might also be applicable for many “planted-versus-planted” problems. Nevertheless, in this paper, we adopt a more conservative view and will explicitly indicate whenever $\mathbb{Q}$ is treated as a planted measure.
The framework in Conjecture 2.2 provides a useful tool for probing the computational feasibility of strong or weak detection. However, as discussed in Section 1.1, it turns out that the failure of strong detection is not enough in our cases, especially when we hope to perform some reductions between statistical models in a regime where weak (but not strong) detection is possible. Thus, in this regime, we aim to characterize a stronger framework that rules out all one-sided test. This motivates thus to propose the following notion of algorithmic contiguity.
Definition 2.3.
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, we say an algorithm $\mathcal{A}$ that takes $\mathsf{Y}$ as input and outputs either 00 or $1$ is a $\mathbb{Q}$-based one-sided test, if
${}\mathbb{Q}(\mathcal{A}(\mathsf{Y})=0)=1-o(1)\mbox{ and }\mathbb{P}(\mathcal{A}(\mathsf{Y})=1)=\Omega(1)\,.$ (2.2)
We say that $\mathbb{P}$ is time-$D$ algorithmic contiguous with respect to $\mathbb{Q}$, denoted as $\mathbb{P}\lhd_{\leq D}\mathbb{Q}$, if no $\mathbb{Q}$-based one-sided testing algorithm runs in time $n^{D}$. We say that $\mathbb{Q}$ and $\mathbb{P}$ are degree-$D$ algorithmic mutually contiguous, denoted as $\mathbb{Q}\bowtie_{\leq D}\mathbb{P}$, if both $\mathbb{Q}\lhd_{\leq D}\mathbb{P}$ and $\mathbb{P}\lhd_{\leq D}\mathbb{Q}$ hold.
Recall that in probability theory we say a sequence of probability measure $\mathbb{P}=\mathbb{P}_{n}$ is contiguous with respect to $\mathbb{Q}=\mathbb{Q}_{n}$, if for all sequence of events $\{A_{n}\}$ we have $\mathbb{Q}_{n}(A_{n})\to 0$ implies that $\mathbb{P}_{n}(A_{n})\to 0$. Thus, our definition can be regarded as the generalization of contiguity in algorithmic view. Our main result in this section can be stated as follows.
Theorem 2.4.
Assuming the Conjecture 2.2, for the high-dimensional hypothesis testing problem between $\mathbb{P}$ and $\mathbb{Q}$, if $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime})=o(1)$ and $\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then we have $\mathbb{P}\lhd_{\leq D/(\log n)^{C}}\mathbb{Q}$ for some constant $C$.
The proof of Theorem 2.4 is incorporated in Section B.1 of the appendix. In the remaining part of this work we will also need a strengthening of Conjecture 2.2, as introduced in Section 1.1. As discussed in the beginning of this section, the low-degree conjecture asserts that (for certain testing problems) low-degree polynomials are at least as powerful as all algorithms of the corresponding runtime (where the correspondence is described in Conjecture 2.2). For our purposes, we introduce a natural refinement of the low-degree conjecture, which posits that low-degree polynomials perform at least as well as all algorithms of the corresponding runtime in terms of the value of the ratio on the right-hand side of (2.1).
Conjecture 2.5 (Revised low-degree conjecture).
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, the following statements hold:
  1. (1)
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$ as $n\to\infty$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then there is a constant $C$ such that there is no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves strong detection between $\mathbb{P}$ and $\mathbb{Q}$.
  2. (2)
    If there is no algorithm with running time $n^{D}$ that achieves strong detection between $\mathbb{P}$ and $\mathbb{Q}$, then for all statistics $f=f(\mathsf{Y})$ that can be computed in running time $n^{D}$, there exists some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ (we allow $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ depend on $f$) such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$ such that $\mathsf{Adv}(f)=O(1)$, where
    $\mathsf{Adv}(f)=\frac{\mathbb{E}_{\mathbb{P}^{\prime}}[f]}{\sqrt{\mathbb{E}_{\mathbb{Q}^{\prime}}[f^{2}]}}\,.$
  3. (3)
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=1+o(1)$ as $n\to\infty$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then there is a constant $C$ such that there is no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves weak detection between $\mathbb{P}$ and $\mathbb{Q}$.
  4. (4)
    If there is no algorithm with running time $n^{D}$ that achieves weak detection between $\mathbb{P}$ and $\mathbb{Q}$, then for all statistics $f=f(\mathsf{Y})$ that can be computed in running time $n^{D}$, there exists some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ (we allow $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ depend on $f$) such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$ such that $\mathsf{Adv}(f)=1+o(1)$).
Note that in Conjecture 2.5 we are allowed to replace $\mathbb{P},\mathbb{Q}$ with some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ that are statistically indistinguishable with $\mathbb{P},\mathbb{Q}$. This modification enables us to avoid some situations where the low-degree advantage explodes due to some “rare events” (see [BAH+22, DDL23+, DMW23+] for example). [^1] In fact, as pointed out by Alexander S. Wein (who learned the following example from Ansh Nagda) in personal communication with the author, such truncation appears necessary for the conjecture to hold. Consider the sparse-PCA problem where the observation $Y=\tfrac{\lambda vv^{\top}}{\|v\|^{2}}+W$ where $W$ is an $n*n$ GOE matrix and $v\in\mathbb{R}^{n}$ having coordinates i.i.d. drawn from the sparse Rademacher prior with parameter $\rho=o(1)$. In this case, when $10<\lambda<$ 1 we have $\mathbb{E}_{\mathbb{P}}[f]/\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}=O(1)$ for all low-degree polynomials $f$ but $\mathbb{E}_{\mathbb{P}}[f]/\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}=\omega(1)$ if $f$ is a thresholding function on $v_{0}^{\top}Yv_{0}$ where $v_{0}$ is a fixed sparse $\{0,1\}$-vector. However, our truncated conjecture still holds as we can take $\mathbb{P}^{\prime}$ to be the conditional measure of $\mathbb{P}$ given $v_{0}^{\top}Yv_{0}$ is small. This revised low-degree conjecture was first proposed in [MW23+] (in a slightly different manner) for a different purpose where they aimed to study the limits of precise error of all computationally feasible algorithms. However, we will show in the next two sections that this conjecture is also useful in performing reductions between different inference tasks.
3 Partial recovery in correlated random graphs
In this section, we will use the framework we established in Section 2 to show the hardness of partial matching in correlated random graphs, thus justifying Items (1) and (2) in Theorem 1.6. To this end, we first state the precise meaning that an algorithm achieves partial matching.
Definition 3.1 (Partial recovery algorithm in correlated Erdős-Rényi model).
Given a sample $(A,B)$ from the law a pair of correlated random graphs in Definition 1.1 (we denote this law of $\mathbb{P}_{*}$). We say an algorithm $\mathcal{A}$ achieves strong partial matching, if it takes $(A,B)$ as input and outputs a family of estimators $\{h_{i,j}:1\leq i,j\leq n\}$ such that
  1. (1)
    $h_{i,j}\in\{0,1\}$ for all $1\leq i,j\leq n$ a.s. under $\mathbb{P}_{*}$;
  2. (2)
    $h_{i,1}+\ldots+h_{i,n}=1$ for all $1\leq i\leq n$ a.s. under $\mathbb{P}_{*}$;
  3. (3)
    There exists a fixed constant $\iota>0$ such that
    $\mathbb{P}_{*}\big{(}h_{1,\pi_{*}(1)}+\ldots+h_{n,\pi_{*}(n)}\geq\iota n\big{)}=1-o(1)\,.$
We say an algorithm $\mathcal{A}$ achieves weak partial matching, if it takes $(A,B)$ as input and outputs a family of estimators $\{h_{i,j}:1\leq i,j\leq n\}$ satisfying Items (1), (2) above and
$\mathbb{P}_{*}\big{(}h_{1,\pi_{*}(1)}+\ldots+h_{n,\pi_{*}(n)}\geq\iota n\big{)}=\Omega(1)\,.$
We point out that in the above definition $h_{i,j}$ can be thought as the estimator of $\mathbf{1}_{\{\pi_{*}(i)=j\}}$, where Item (3) implies this algorithm correctly matches a positive fraction of $\pi_{*}(i)$ with probability $1-o(1)$ (or with positive probability) and Items (1) and (2) are some regularity requirements. Our result in this section can be stated as follows:
Theorem 3.2.
Assuming Conjecture 2.5, we have the following:
  1. (1)
    For the correlated Erdős-Rényi graphs $\mathcal{G}(n,q,\rho)$ where $q=n^{-1+o(1)}$ and $\rho<\sqrt{\alpha}-\delta$ for a fixed constant $\delta>0$. There exists a constant $C$ such that no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves weak partial matching, provided that
    ${}D=\exp\Big{(}o\big{(}\tfrac{\log n}{\log(nq)}\wedge\sqrt{\log n}\big{)}\Big{)}\,.$ (3.1)
  2. (2)
    For the correlated SBMs $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ where $\lambda=O(1)$ and $\epsilon^{2}\lambda<1-\delta,s<\sqrt{\alpha}-\delta$ for a fixed constant $\delta>0$. There exists a constant $C$ such that no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves strong partial matching, provided that
    ${}D=n^{o(1)}\,.$ (3.2)
The main step of proving Theorem 3.2 is to show the following proposition:
Proposition 3.3.
Assuming Conjecture 2.5, we have the following:
  1. (1)
    If $(G,A,B)\sim\mathcal{G}(n,q,\rho)$ and let $\mathbb{P}_{*}$ to be the joint law of $(\pi_{*},G,A,B)$ where $\pi_{*}$ is the latent matching. Suppose that $q,\rho,D$ satisfy the assumptions in Item (1) of Theorem 3.2. Then there exists a constant $C$ such that for all $\{f_{i,j}:1\leq i,j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$, we have $\mathbb{E}_{\mathbb{P}_{*}}[f_{i,\pi_{*}(i)}]=o(1)$ for all $1\leq i\leq n$.
  2. (1)
    If $(G,A,B)\sim\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon,s)$ and let $\mathbb{P}_{*}$ to be the joint low of $(\pi_{*},G,A,B)$ where $\pi_{*}$ is the latent matching. Suppose that $\lambda,k,\epsilon,s$ and $D$ satisfy the assumptions in Item (2) of Theorem 3.2. Then there exists a constant $C$ and an event $\mathcal{E}$ such that $\mathbb{P}_{*}(\mathcal{E})=\Omega(1)$, and for all $\{f_{i,j}:1\leq i,j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ we have $\mathbb{E}_{\mathbb{P}_{*}}[f_{i,\pi_{*}(i)}\mid\mathcal{E}]=o(1)$ for all $1\leq i\leq n$.
Clearly, based on Proposition 3.3, we can deduce Theorem 3.2 via a simple Markov inequality. The rest of this section is devoted to the proof of Proposition 3.3. In the following subsections, our main focus is on proving Item (1) of Proposition 3.3. Given the similarity between the proofs of Item (1) and Item (2), for Item (2) we will provide an outline with the main differences while adapting arguments from proving Item (1) without presenting full details. [^2] After the completion of an earlier version of this draft, we realized that actually we can prove Proposition 3.3 (and thus proving Theorem 3.3) assuming only a truncated version of Conjecture 2.2 (which is weaker than Conjecture 2.5). However, this alternative proof seems to be more problem dependent so we defer it to Section C of the appendix. In the main text, we focus on proving Proposition 3.3 under the stronger assumption of Conjecture 2.5.
3.1 Proof of Item (1) in Proposition 3.3
This subsection is devoted to the proof of Item (1) in Proposition 3.3. Throughout this subsection, we will denote $\mathbb{P}_{*}$ to be the law of $(\pi_{*},G,A,B)$ where $(G,A,B)\sim\mathcal{G}(n,q;\rho)$. We will also denote $\mathbb{P}$ to be the marginal law of $(A,B)$. In addition, we assume throughout this subsection that there exists a small constant $0.010<\delta<$ 0.01 such that
${}\rho^{2}<\alpha-\delta\,,\quad q=n^{-1+o(1)}\,,\quad\log D=o\Big{(}\tfrac{\log n}{\log(nq)}\wedge\sqrt{\log n}\Big{)}\,.$ (3.3)
We first introduce some notations used in [DDL23+].
Definition 3.4.
Given a graph $H=H(V,E)$, define
$\Phi(H)={\big{(}n^{1+4/D}{D}^{20}\big{)}^{|V(H)|}\big{(}qD^{6}\big{)}^{|E(H)|}\,,}$ (3.4)
and the graph $H$ is said to be bad if ${\Phi(H)<(\log n)^{-1}}$. Furthermore, we say a graph is admissible if it contains no bad subgraph, and we say it is inadmissible otherwise.
Denote $\mathcal{E}$ for the event that $G$ does not contain any bad subgraph with no more than $d^{2}$ vertices. In addition, let $\overline{\mathbb{P}}_{*}$ be the conditional version of $\mathbb{P}_{*}$ given $\mathcal{E}$, and let $\overline{\mathbb{P}}$ be the corresponding marginal distribution of $\overline{\mathbb{P}}_{*}$ on $(A,B)$.
We remark here that our definition of “bad” amounts to an atypically large edge density, with a carefully chosen quantitative threshold on “large”. Roughly speaking, we expect that any subgraph with size no more than $D^{2}=n^{o(1)}$ of a sparse Erdős-Rényi graph has edge-to-vertex ratio $1+o(1)$. In the definition of $\Phi$, the term $n^{1+4/D}D^{20}$ should be thought as $n^{1+o(1)}$, and $qD^{6}$ as $n^{-1+o(1)}$. The $o(1)$ terms are tuned carefully so that for a typical subgraph $H$ of a sparse Erdős-Rényi graph, $\Phi(H)$ is much bigger than $1$. The choice of $(\log n)^{-1}$ as the $\Phi$-threshold for bad graph is somewhat arbitrary, which we will only need to be vanishing as $n\to\infty$. In [DDL23+], the authors showed that one the one hand, we have $\mathbb{P}_{*}(\mathcal{E})=1-o(1)$ and thus $\operatorname{TV}(\mathbb{P}_{*},\overline{\mathbb{P}}_{*}),\operatorname{TV}(\mathbb{P},\overline{\mathbb{P}})=o(1)$; on the other hand, we have $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta}(1)$, thus verifying the low-degree hardness for the detection problem. The first step of our proof is to generalize the result in [DDL23+], as incorporated in the following lemma.
Lemma 3.5.
For all $1\leq i,j\leq n$, we have $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}(\cdot\mid\pi_{*}(i)=j)}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta}(1)$.
The proof of Lemma 3.5 is quite technical and thus we postpone it to Section B.2 of the appendix. Now, based on Lemma 3.5, we establish the following result, which basically suggests that it is impossible to obtain a “good approximation” of $\mathbf{1}_{\{\pi_{*}(i)=j\}}$ in some sense.
Lemma 3.6.
Assuming Conjecture 2.5, there exists constant $C$ such that for all $1\leq i\leq n$ and all statistics $\{g_{i,j}=g_{i,j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$, it holds that
${}\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\Big{[}\big{(}\mathbf{1}_{\{\pi_{*}(i)=j\}}-g_{i,j}\big{)}^{2}\Big{]}\geq 1-o(1)\,.$ (3.5)
Proof.
Without losing of generality, we may assume that $i=1$ in the following proof. Suppose on the contrary that there are statistics $\{g_{j}=g_{j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ such that
${}\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\Big{[}\big{(}\mathbf{1}_{\{\pi_{*}(1)=j\}}-g_{j}\big{)}^{2}\Big{]}\leq 1-c\mbox{ for some constant }c>0\,.$ (3.6)
Without losing of generality, we may assume that $0\leq g_{i}\leq 1$, since otherwise we may replace $g_{i}$ with $\min\{\max\{g_{i},0\},1\}$, which will only make the left hand side of (3.6) smaller. Denote
${}\Lambda:=\Big{\{}1\leq j\leq n:\mathbb{E}_{\mathbb{P}_{*}}\big{[}(\mathbf{1}_{\{\pi_{*}(1)=j\}}-g_{j})^{2}\big{]}\leq\frac{1-\frac{c}{2}}{n}\Big{\}}\,.$ (3.7)
Using Markov inequality, we see that
${}|\Lambda|\geq n-\frac{1-c}{1-\frac{c}{2}}n\geq\frac{cn}{2}\,.$ (3.8)
Recall that $\overline{\mathbb{P}}=\mathbb{P}(\cdot\mid\mathcal{E})$ where $\mathcal{E}$ is independent to the latent matching $\pi_{*}$ and $\mathbb{P}_{*}(\mathcal{E})=1-o(1)$. Thus, it is straightforward to check that for all $i\in\Lambda$
$\mathbb{E}_{\overline{\mathbb{P}}}\big{[}(1-g_{i})^{2}\mid\pi_{*}(1)=i\big{]}\leq[1+o(1)]n\cdot\mathbb{E}_{\mathbb{P}_{*}}\big{[}(\mathbf{1}_{\{\pi_{*}(1)=i\}}-g_{i})^{2}\big{]}\leq 1-\tfrac{c}{2}\,.$
Thus, using Markov inequality we see that (note that $0\leq g_{i}\leq 1$)
¯$\overline{\mathbb{P}}(g_{i}>\tfrac{c}{2}\mid\pi_{*}(1)=i)\geq\Omega(1)\,.$ (3.9)
However, using Lemma 3.5 and Theorem 2.4, we see that (assuming Conjecture 2.5) we have
¯$\overline{\mathbb{P}}(\cdot\mid\pi_{*}(1)=i)\lhd_{\leq D}\mathbb{Q}\,.$
Thus we must have $\mathbb{Q}(g_{i}>\tfrac{c}{2})\geq\Omega(1)$. This yields that
${}\mathbb{Q}^{\prime}(g_{i}>\tfrac{c}{2})\geq\Omega(1)\mbox{ for all }\operatorname{TV}(\mathbb{Q}^{\prime},\mathbb{Q})=o(1)\,.$ (3.10)
This further suggests that
${}\mathbb{E}_{\mathbb{Q}^{\prime}}[g_{i}]=\Omega(1)\mbox{ for all }\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)\mbox{ and }i\in\Lambda\,.$ (3.11)
To this end, define $g=g_{1}+\ldots+g_{n}$. Using (3.8) and (3.11), we see that (recall again that $0\leq g_{i}\leq 1$)
${}\mathbb{E}_{\mathbb{Q}^{\prime}}[g]=\Omega(n)\mbox{ for all }\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)\,.$ (3.12)
However, using $0\leq g\leq n$, we see that for all $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime})=o(1)$ we have
$\mathbb{E}_{\mathbb{P}^{\prime}}[g^{2}]$ $\leq\operatorname{TV}(\overline{\mathbb{P}},\mathbb{P}^{\prime})\cdot\|g\|_{\infty}^{2}+\mathbb{E}_{\overline{\mathbb{P}}}[g^{2}]\leq o(n^{2})+2\big{(}1+\mathbb{E}_{\overline{\mathbb{P}}}[(g-1)^{2}]\big{)}$
$\leq o(n^{2})+2\mathbb{E}_{\overline{\mathbb{P}}}\Big{[}\Big{(}\sum_{i=1}^{n}(\mathbf{1}_{\{\pi_{*}(1)=i\}}-g_{i})\Big{)}^{2}\Big{]}$
[…]
Combining (3.12) and (3.13), we see that for all $\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime}),\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime})=o(1)$ we have
${}\frac{\mathbb{E}_{\mathbb{Q}^{\prime}}[g]}{\sqrt{\mathbb{E}_{\mathbb{P}^{\prime}}[g^{2}]}}=\omega(1)\,.$ (3.14)
Thus, using Item (2) in Conjecture 2.5 on the testing problem between $\mathbb{P}$ and $\mathbb{Q}$, we see that there exists an algorithm with running time $n^{D/(\log n)^{C}}$ that strongly distinguish $\mathbb{P}$ and $\mathbb{Q}$. However, it is impossible assuming Item (1) of Conjecture 2.5 as [DDL23+] showed that $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta}(1)$ where $\operatorname{TV}(\overline{\mathbb{P}},\mathbb{P})=o(1)$. This implies that (3.6) is impossible and thus completes our proof. ∎
Now we can finish the proof of Item (1) in Proposition 3.3.
Proof of Item (1) in Proposition 3.3.
We choose $C$ as in the statement of Lemma 3.6. Suppose on the contrary there are statistics $\{f_{j}:1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ with $\mathbb{E}_{\mathbb{P}_{*}}[f_{\pi_{*}(i)}]\geq 1-c$ for some fixed constant $0.010<c<$ 0.01. Using Items (1) and (2) in Definition 3.1, we see that $f_{j}f_{k}=0$ for all $k\neq j$, and thus
$1=\mathbb{E}_{\mathbb{P}}\big{[}(f_{1}+\ldots+f_{n})^{2}\big{]}=\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}}\big{[}f_{j}^{2}\big{]}\,.$ (3.15)
In addition, we have
$\mathbb{E}_{\mathbb{P}_{*}}\big{[}f_{\pi_{*}(i)}\big{]}=\frac{1}{n}\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\big{[}f_{j}\mid\pi_{*}(i)=j\big{]}\geq c\,.$ (3.16)
Thus, for all $\lambda\in[0,1]$ we have
[…]
Thus, by choosing $\lambda=\lambda(c)$ to be a sufficiently small positive constant we get that
$\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\Big{[}\big{(}\mathbf{1}_{\{\pi_{*}(i)=j\}}-g_{j}\big{)}^{2}\Big{]}=1-\Omega(1)\mbox{ where }g_{j}=\tfrac{1-\lambda}{n}+\lambda f_{j}\,,$
contradicting to Lemma 3.6. This leads to the desired result. ∎
3.2 Proof of Item (2) in Proposition 3.3
This subsection is devoted to the proof of Item (2) in Proposition 3.3. Recall Definitions 1.3 and 1.4. Throughout this subsection, we will denote $\mathbb{P}_{*}$ to be the joint law of $(\pi_{*},\sigma_{*},G,A,B)$ where $(G,A,B)\sim\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ and $\mathbb{P}$ the marginal law of $(A,B)$. In addition, we denote $\mathbb{Q}$ to be the law of a pair of independent Erdős-Rényi models $\mathcal{G}(n,\tfrac{\lambda s}{n})$. In addition, we assume throughout this subsection that there exists a small constant $0.010<\delta<$ 0.01 such that
${}s<\sqrt{\alpha}-\delta\,,\quad\epsilon^{2}\lambda s<1-\delta\,.$ (3.17)
We also choose a sufficiently large constant $N=N(k,\lambda,\delta,\epsilon,s)\geq 2/\delta$ such that
$(\sqrt{\alpha}-\delta)(1+\epsilon^{N}k)\leq\sqrt{\alpha}-\delta/2\,;\quad 10k(1-\delta)^{N}\leq(1-\delta/2)^{N}\,;$ (3.18)
$(\sqrt{\alpha}-\delta/2)(1+(1-\delta/2)^{N})^{2}\leq\sqrt{\alpha}-\delta/4\,;\quad(1-\delta/2)^{N}(N+1)\leq 1\,.$
We first show how to construct the event $\mathcal{E}$ in Item (2) in Proposition 3.3.
Definition 3.7.
Denote $\tilde{\lambda}=\lambda\vee 1$. Given a graph $H=H(V,E)$, define
$\Upsilon(H)=\Big{(}\frac{2\tilde{\lambda}^{2}k^{2}n}{D^{50}}\Big{)}^{|V(H)|}\Big{(}\frac{1000\tilde{\lambda}^{20}k^{20}D^{50}}{n}\Big{)}^{|E(H)|}\,.$ (3.19)
Then we say the graph $H$ is bad if $\Upsilon(H)<(\log n)^{-1}$, and we say a graph $H$ is self-bad if $H$ is bad and $\Upsilon(H)<\Upsilon(K)$ for all $K\subset H$. Furthermore, we say that a graph $H$ is admissible if it contains no bad subgraph and $\mathtt{C}_{j}(H)=\emptyset$ for $j\leq N$; we say $H$ is inadmissible otherwise. Denote $\mathcal{E}=\mathcal{E}^{(1)}\cap\mathcal{E}^{(2)}$, where $\mathcal{E}^{(1)}$ is the event that $G$ does not contain any bad subgraph with no more than $D^{3}$ vertices, and $\mathcal{E}^{(2)}$ is the event that $G$ does not contain any cycles with length at most $N$.
Definition 3.8.
List all self-bad subgraphs of $\mathcal{K}_{n}$ with at most $D^{3}$ vertices and all cycles of $\mathcal{K}_{n}$ with lengths at most $N$ in an arbitrary but prefixed order $(B_{1},\ldots,B_{\mathtt{M}})$. Define a stochastic block model with “bad graphs” removed as follows: (1) sample $G\sim\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon)$; (2) for each $\mathtt{1}\leq\mathtt{i}\leq\mathtt{M}$ such that $B_{\mathtt{i}}\subset G$, we independently uniformly remove one edge in $B_{\mathtt{i}}$. The unremoved edges in $G$ constitute a graph $G^{\prime}$, which is the output of our modified stochastic block model. Clearly, from this definition $G^{\prime}$ does not contain any cycle of length at most $N$ nor any bad subgraph with at most $D^{3}$ vertices. Conditioned on $G^{\prime}$ and $\pi_{*}$, we define
$A^{\prime}_{i,j}=G^{\prime}_{i,j}J^{\prime}_{i,j},B^{\prime}_{i,j}=G^{\prime}_{\pi_{*}^{-1}(i),\pi_{*}^{-1}(j)}K^{\prime}_{i,j}\,,$
where $J^{\prime}$ and $K^{\prime}$ are independent Bernoulli variables with parameter $s$. Let ~=~$\widetilde{\mathbb{P}}_{*}=\widetilde{\mathbb{P}}_{*,n}$ be the law of $(\sigma_{*},\pi_{*},G,G^{\prime},A^{\prime},B^{\prime})$ and denote ~=~$\widetilde{\mathbb{P}}=\widetilde{\mathbb{P}}_{n}$ the marginal law of $(A^{\prime},B^{\prime})$.
It was shown in [CDGL24+, Lemmas 4.2 and 4.4] that
$\mathbb{P}_{*}(\mathcal{E})=\Omega(1)\mbox{ and }\operatorname{TV}(\widetilde{\mathbb{P}},\mathbb{P}(\cdot\mid\mathcal{E}))=o(1)\,.$
Similarly as in Section 3.1, our first step is to show the following lemma.
Lemma 3.9.
We have $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\widetilde{\mathbb{P}}(\cdot\mid\pi_{*}(i)=j)}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta,k}(1)$.
The proof of Lemma 3.9 is incorporated in Section B.3 of the appendix. Based on Lemma 3.9, we can deduce our main result just as how we deduce Theorem 3.2 from Lemma 3.5. The only difference is that we will replace all $\overline{\mathbb{P}}$ with $\widetilde{\mathbb{P}}$ and replace all $\mathbb{P}$ with $\mathbb{P}(\cdot\mid\mathcal{E})$ so we omit further details here.
4 Detection in correlated SBMs
In this section, we will use the framework we established in Section 2 in stochastic block models below KS-threshold. The main results of this section is incorporated as follows.
Theorem 4.1.
For any constant $K\in\mathbb{N}$, denote $\mathbb{P}$ to be the law of $K$ independent stochastic block models $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon)$ and denote $\mathbb{Q}$ to be the law of $K$ independent Erdős-Rényi graphs $\mathcal{G}(n,\tfrac{\lambda}{n})$. Then, assuming Conjecture 2.2, for any $\delta>0$ there exists $\lambda_{0}=\lambda_{0}(\delta,k)$ to be a sufficiently large constant such that when $\epsilon^{2}\lambda<1-\delta$ and $\lambda>\lambda_{0}$, we have we have $\mathbb{P}\bowtie_{\leq D}\mathbb{Q}$ for any $D=n^{o(1)}$.
Our result has an immediate corollary in the detection problem between a pair of correlated SBMs and a pair of independent SBMs, as incorporated in the following corollary.
Corollary 4.2.
Assuming Conjecture 2.2, when $\epsilon^{2}\lambda s<1-\delta$, $s<\sqrt{\alpha}-\delta$ and $\lambda>\lambda_{0}(\delta,k)$ there is no algorithms with polynomial running time that can strongly distinguish $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ and two independent $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$.
Proof.
Denote $\mu_{\mathsf{CorSBM}},\mu_{\mathsf{IndSBM}}$ and $\mu_{\mathsf{IndER}}$ to be the law of a pair of correlated SBMs $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$, a pair of independent SBMs $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$ and a pair of independent Erdős-Rényi graphs $\mathcal{G}(n,\tfrac{\lambda s}{n})$, respectively. Suppose on the contrary that there exists an efficient algorithm $\mathcal{A}$ such that
${}\mu_{\mathsf{CorSBM}}\big{(}\mathcal{A}(A,B)=1\big{)}=1-o(1)\,,\quad\mu_{\mathsf{IndSBM}}\big{(}\mathcal{A}(A,B)=0\big{)}=1-o(1)\,.$ (4.1)
Using Theorem 4.1 with $K=2$, we see that $\mu_{\mathsf{IndSBM}}\bowtie_{\leq D}\mu_{\mathsf{IndER}}$ for $D=n^{o(1)}$. Thus
$\mu_{\mathsf{IndER}}\big{(}\mathcal{A}(A,B)=0\big{)}=1-o(1)\,.$
Thus, this algorithm $\mathcal{A}$ strongly distinguish $\mu_{\mathsf{CorSBM}}$ and $\mu_{\mathsf{IndER}}$. This contradicts Conjecture 2.2 and the low-degree hardness established in [CDGL24+, Theorem 1.3]. ∎
The rest part of this section is devoted to the proof of Theorem 4.1. For notational simplicity, in the following we will only prove the case where $K=1$ and the proof for general $K$ is similar. Note that $\epsilon^{2}\lambda<1-\delta$ and $\lambda>\lambda_{0}$ implies that $\epsilon<\epsilon_{0}=\lambda_{0}^{-1/2}$. We choose $\lambda_{0}=\lambda_{0}(\delta,k)$ to be a sufficient large constant such that
${}(1-\delta)^{-1}\cdot\tfrac{(k-1)\sqrt{1-\epsilon}+\sqrt{1+\epsilon(k-1)}}{k}\geq(1-\delta/2)^{-1}\mbox{ for all }\epsilon<\epsilon_{0}=\lambda_{0}^{-1/2}\,.$ (4.2)
In the rest part of this section we will always assume that
${}D=n^{o(1)}\,,\lambda>\lambda_{0}\mbox{ and }\epsilon^{2}\lambda<1-\delta\mbox{ for some constant }0<\delta<0.01\,.$ (4.3)
Clearly, using Theorem 2.4, it suffices to show that under (4.3) and $\lambda_{0}$ we have
$\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta,k}(1)\mbox{ and }\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\big{)}=O_{\delta,k}(1)\,.$
Indeed, it has been shown in [HS17] that $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\big{)}=O_{\delta,k}(1)$ provided with (4.3). It remains to show that under (4.3) we have (note that now $\mathbb{P}$ is the planted measure)
${}\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\big{)}=\sup_{f\in\mathcal{P}_{D}}\frac{\mathbb{E}_{\mathbb{Q}}[f]}{\sqrt{\mathbb{E}_{\mathbb{P}}[f^{2}]}}=O_{\delta,k}(1)\,.$ (4.4)
We point out that our approach to proving (4.4) is based on the the work [SW25]. To this end, define
${}\omega(\sigma_{i},\sigma_{j})=\begin{cases}k-1\,,&\sigma_{i}=\sigma_{j}\,;\\ -1\,,&\sigma_{i}\neq\sigma_{j}\end{cases}$ (4.5)
In addition, for all $S\Subset\mathcal{K}_{n}$ define
${}\phi_{S}\big{(}\{G_{i,j}\}\big{)}=\prod_{(i,j)\in E(S)}\frac{G_{i,j}-\frac{\lambda}{n}}{\sqrt{\frac{\lambda}{n}(1-\frac{\lambda}{n})}}\,.$ (4.6)
It is well known in [HS17] that $\{\phi_{S}:S\Subset\mathcal{K}_{n},|E(S)|\leq D\}$ constitutes a standard orthogonal basis of $\mathcal{P}_{D}$ under $\mathbb{Q}$. Thus, each $f\in\mathcal{P}_{D}$ can be written as
${}f(G)=\sum_{S\Subset\mathcal{K}_{n},|E(S)|\leq D}\widehat{f}_{S}\cdot\phi_{S}(G)\,,$ (4.7)
which means that $f$ is uniquely characterized by a vector $\widehat{f}$ indexed by $\{S\Subset\mathcal{K}_{n}:|E(S)|\leq D\}$. In addition, direct calculation yields that
${}\mathbb{E}_{\mathbb{Q}}\big{[}\phi_{S}(G)\big{]}=\mathbf{1}_{\{S=\emptyset\}}\,.$ (4.8)
Thus, we have
[…]
where $c$ is a vector indexed by $\{S\Subset\mathcal{K}_{n}:|E(S)|\leq D\}$ with
${}c_{S}=\mathbf{1}_{\{S=\emptyset\}}\,.$ (4.10)
We now turn to $\mathbb{E}_{\mathbb{P}}[f^{2}]$. For any $\sigma\in[k]^{n}$ and $S\Subset\mathcal{K}_{n}$, define
${}\psi_{\sigma,S}\big{(}\{G_{i,j}\}\big{)}=k^{\frac{n}{2}}\mathbf{1}_{\sigma_{*}=\sigma}\cdot\prod_{(i,j)\in E(S)}\frac{G_{i,j}-\frac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n}}{\sqrt{\frac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n}(1-\frac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n})}}$ (4.11)
We can check that $\{\psi_{\sigma,S}:\sigma\in[k]^{n},S\Subset\mathcal{K}_{n}\}$ is standard orthogonal under $\mathbb{P}_{*}$, i.e., we have
${}\mathbb{E}_{\mathbb{P}_{*}}\big{[}\psi_{\sigma,S}\psi_{\sigma^{\prime},S^{\prime}}\big{]}=\mathbf{1}_{\{\sigma=\sigma^{\prime},S=S^{\prime}\}}\,.$ (4.12)
Lemma 4.3.
We have
${}\mathbb{E}_{\mathbb{P}}\big{[}\phi_{S}(G)\psi_{\sigma,H}(G)\big{]}=\frac{\mathbf{1}_{H\subset S}}{k^{\frac{n}{2}}}\prod_{(i,j)\in E(H)}\mathtt{h}(\sigma_{i},\sigma_{j})\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\sqrt{\tfrac{\epsilon^{2}\lambda}{n}}\,,$ (4.13)
where
${}\mathtt{h}(\sigma_{i},\sigma_{j})=\sqrt{\frac{(1-\frac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n})(1+\epsilon\omega(\sigma_{i},\sigma_{j}))}{1-\frac{\lambda}{n}}}\,.$ (4.14)
Proof.
Note that using (4.6) and (4.11), we have that $\mathbb{E}_{\mathbb{P}}[\phi_{S}(G)\psi_{\sigma,H}(G)]$ equals
$\mathbb{E}_{\mathbb{P}}\Bigg{[}k^{\frac{n}{2}}\mathbf{1}_{\sigma_{*}=\sigma}\cdot\prod_{(i,j)\in E(S)}\frac{G_{i,j}-\frac{\lambda}{n}}{\sqrt{\frac{\lambda}{n}(1-\frac{\lambda}{n})}}\cdot\prod_{(i,j)\in E(H)}\frac{G_{i,j}-\frac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n}}{\sqrt{\frac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n}(1-\frac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n})}}\Bigg{]}$
=\ = $k^{-\frac{n}{2}}\mathbb{E}_{\mathbb{P}_{\sigma}}\Bigg{[}\prod_{(i,j)\in E(S)}\frac{G_{i,j}-\frac{\lambda}{n}}{\sqrt{\frac{\lambda}{n}(1-\frac{\lambda}{n})}}\cdot\prod_{(i,j)\in E(H)}\frac{G_{i,j}-\frac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n}}{\sqrt{\frac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n}(1-\frac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n})}}\Bigg{]}\,.$ (4.15)
Clearly, when $H\not\subset S$ we have (4.15) cancels to 00, since (denote $(i,j)$ to be an element in $E(H)\setminus E(S)$) under $\mathbb{P}_{\sigma}$ we have $\mathbb{E}_{\mathbb{P}_{\sigma}}[G_{i,j}-\tfrac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n}]=0$ and all the edges $\{G_{i,j}:(i,j)\in\operatorname{U}_{n}\}$ are independent. In addition, from direct calculation we have that
$\mathbb{E}_{\mathbb{P}_{\sigma}}\big{[}G_{i,j}-\tfrac{\lambda}{n}\big{]}=\tfrac{\epsilon\lambda\omega(\sigma_{i},\sigma_{j})}{n}\,,$
$\mathbb{E}_{\mathbb{P}_{\sigma}}\big{[}(G_{i,j}-\tfrac{\lambda}{n})(G_{i,j}-\tfrac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n})\big{]}=\tfrac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n}\big{(}1-\tfrac{(1+\epsilon\omega(\sigma_{i},\sigma_{j}))\lambda}{n}\big{)}\,.$
Plugging this result into (4.15), we get (4.13) and thus completes the proof. ∎
Based on Lemma 4.3, we define a matrix $M$ with rows indexed by $\{S:S\Subset\mathcal{K}_{n},|E(S)|\leq D\}$ and columns indexed by $\{(\sigma,S):\sigma\in[k]^{n},S\Subset\mathcal{K}_{n},|E(S)|\leq D\}$, and entries given by
${}M_{S;(\sigma,H)}=\frac{\mathbf{1}_{H\subset S}}{k^{\frac{n}{2}}}\prod_{(i,j)\in E(H)}\mathtt{h}(\sigma_{i},\sigma_{j})\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\sqrt{\tfrac{\epsilon^{2}\lambda}{n}}\,.$ (4.16)
From Parserval’s inequality, we see that
[…]
$=\sum_{\begin{subarray}{c}\sigma\in[k]^{n},H\Subset\mathcal{K}_{n}\\ |E(H)|\leq D\end{subarray}}\Big{(}\sum_{\begin{subarray}{c}S\Subset\mathcal{K}_{n}\\ |E(S)|\leq D\end{subarray}}\widehat{f}_{S}M_{S;(\sigma,H)}\Big{)}^{2}=\big{\|}\widehat{f}M\big{\|}^{2}\,.$ (4.17)
Thus, we have
${}\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}\big{)}=\sup_{f\in\mathcal{P}_{D}}\Bigg{\{}\frac{\mathbb{E}_{\mathbb{Q}}[f]}{\sqrt{\mathbb{E}_{\mathbb{P}}[f^{2}]}}\Bigg{\}}\leq\sup_{\widehat{f}}\Bigg{\{}\frac{\langle\widehat{f},c\rangle}{\|\widehat{f}M\|}\Bigg{\}}\leq\inf_{Mu^{\top}=c}\big{\{}\|u\|\big{\}}\,,$ (4.18)
where the last inequality follows from the fact that for $Mu_{0}^{\top}=c$ we have
$\langle\widehat{f},c\rangle=\langle\widehat{f},Mu_{0}^{\top}\rangle=\langle\widehat{f}M,u_{0}\rangle\leq\|u_{0}\|\cdot\|\widehat{f}M\|\,.$
Regarding (4.18), it suffices to show that there exists $Mu^{\top}=c$ and $\|u\|=O_{\delta,k}(1)$. Note that $Mu^{\top}=c$ is equivalent to
${}\sum_{\sigma\in[k]^{n}}\sum_{H\subset S}\frac{u_{\sigma,H}}{k^{\frac{n}{2}}}\prod_{(i,j)\in E(H)}\mathtt{h}(\sigma_{i},\sigma_{j})\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\sqrt{\tfrac{\epsilon^{2}\lambda}{n}}=\mathbf{1}_{S=\emptyset}\,.$ (4.19)
The first step of our analysis is to simplify (4.19). Based on Lemma A.7, we see that for all $H\subset S$ with $\mathcal{L}(S)\setminus V(H)\neq\emptyset$, we have (denote $\mathtt{V}=V(H)$)
$\sum_{\sigma\in[k]^{n}}u_{\sigma,H}\prod_{(i,j)\in E(H)}\mathtt{h}(\sigma_{i},\sigma_{j})\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\sqrt{\tfrac{\epsilon^{2}\lambda}{n}}$
=\ = $\big{(}\tfrac{\epsilon^{2}\lambda}{n}\big{)}^{\frac{|E(S)|-|E(H)|}{2}}\sum_{\sigma_{\mathtt{V}}\in[n]^{\mathtt{V}}}\prod_{(i,j)\in E(H)}\mathtt{h}(\sigma_{i},\sigma_{j})\times\sum_{\sigma_{[n]\setminus\mathtt{V}}\in[k]^{[n]\setminus\mathtt{V}}}\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})=0\,.$
Thus, (4.19) can be further simplified to
${}\sum_{\sigma\in[k]^{n}}\sum_{\begin{subarray}{c}H\subset S\\ \mathcal{L}(S)\subset V(H)\end{subarray}}\frac{u_{\sigma,H}}{k^{\frac{n}{2}}}\prod_{(i,j)\in E(H)}\mathtt{h}(\sigma_{i},\sigma_{j})\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\sqrt{\tfrac{\epsilon^{2}\lambda}{n}}=\mathbf{1}_{S=\emptyset}\,.$ (4.20)
We now construct the solution $\{u_{\sigma,H}:\sigma\in[k]^{n},H\Subset\mathcal{K}_{n},|E(H)|\leq D\}$ of (4.20) as follows: let $u_{\sigma,H}=\frac{1}{k^{\frac{n}{2}}}\cdot\Xi(H)$, where
Ξ()=1 and ${}\Xi(\emptyset)=1\mbox{ and }\Xi(S)=0\mbox{ for }\mathcal{L}(S)\neq\emptyset$ (4.21)
and then iteratively define for all $\mathcal{L}(S)=\emptyset$ by
$\Xi(S)=$ $-\Big{(}\mathbb{E}_{\sigma\sim\nu}\Big{[}\prod_{(i,j)\in E(S)}\mathtt{h}(\sigma_{i},\sigma_{j})\Big{]}\Big{)}^{-1}\sum_{\begin{subarray}{c}H\subset S\\ \mathcal{L}(H)=\emptyset\end{subarray}}\big{(}\tfrac{\epsilon^{2}\lambda}{n}\big{)}^{\frac{|E(S)|-|E(H)|}{2}}\Xi(H)$ (4.22)
×$\times\ \mathbb{E}_{\sigma\sim\nu}\Big{[}\prod_{(i,j)\in E(H)}\mathtt{h}(\sigma_{i},\sigma_{j})\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\Big{]}\,,$
To prove Theorem 4.1, it suffices to show the following lemma, as incorporated in Section B.4 of the appendix.
Lemma 4.4.
The vector $\{u_{\sigma,H}:\sigma\in[k]^{n},H\Subset\mathcal{K}_{n},|E(H)|\leq D\}$ satisfies (4.20) and $\|u\|=O_{\delta,k}(1)$.
Acknowledgment. The author thanks Jian Ding and Jingqiu Ding for helpful discussions, Alexander S. Wein for helpful comments on the revised low-degree conjecture, and Hang Du for pointing him to the reference [SW25]. This work is partially supported by National Key R&\&&D program of China (No. 2023YFA1010103) and NSFC Key Program (Project No. 12231002).
Appendix A Preliminary results in graphs and probability
In this section we include some preliminary results we will use in the main part of the paper. We first state some results on the structural properties of graphs established in [DDL23+, CDGL24+].
Lemma A.1 ([DDL23+], Lemma A.1).
Let $S,T\Subset\mathcal{K}_{n}$ satisfy $S\cong\mathbf{S}$ and $T\cong\mathbf{T}$ for some $\mathbf{S},\mathbf{T}\in\mathcal{H}$. Recall $S\Cup T,S\Cap T\Subset\mathcal{K}_{n}$ defined as edge-induced graphs of $\mathcal{K}_{n}$. We have the following hold:
  1. (i)
    $|V(S\cup T)|+|V(S\Cap T)|\leq|V(S)|+|V(T)|$, and $|E(S\cup T)|+|E(S\Cap T)|=|E(S)|+|E(T)|$.
  2. (ii)
    $\Phi(S\cup T)\Phi(S\Cap T)\leq\Phi(S)\Phi(T)$.
  3. (iii)
    If $\mathbf{S}\subset\mathbf{T}$, then $|\!\operatorname{Aut}(\mathbf{S})|\leq|\!\operatorname{Aut}(\mathbf{T})|\cdot|V(\mathbf{T})|^{2(|E(\mathbf{T})|-|E(\mathbf{S})|)}$.
  4. (iv)
    #$\#\big{\{}T^{\prime}\Subset\mathcal{K}_{n}:S\Subset T^{\prime},|V(T^{\prime})|-|V(S)|=k,|E(T^{\prime})|-|E(S)|=l\big{\}}\leq n^{k}(|V(S)|+k)^{2l}$.
  5. (v)
    #$\#\big{\{}T^{\prime}\Subset S:|E(S)|-|E(T^{\prime})|=k\big{\}}\leq|E(S)|^{k}$.
Lemma A.2 ([CDGL24+], Lemma A.3).
For $H\subset S$, we can decompose $E(S)\setminus E(H)$ into $\mathtt{m}$ cycles ${C}_{\mathtt{1}},\ldots,{C}_{\mathtt{m}}$ and $\mathtt{t}$ paths ${P}_{\mathtt{1}},\ldots,{P}_{\mathtt{t}}$ for some $\mathtt{m},\mathtt{t}\geq 0$ such that the following holds:
  1. (1)
    ${C}_{\mathtt{1}},\ldots,{C}_{\mathtt{m}}$ are vertex-disjoint (i.e., $V(C_{\mathtt{i}})\cap V(C_{\mathtt{j}})=\emptyset$ for all $\mathtt{i}\neq\mathtt{j}$) and $V(C_{\mathtt{i}})\cap V(H)=\emptyset$ for all $1\leq\mathtt{i}\leq\mathtt{m}$.
  2. (2)
    $\operatorname{EndP}({P}_{\mathtt{j}})\subset V(H)\cup(\cup_{\mathtt{i}=1}^{\mathtt{m}}V(C_{\mathtt{i}}))\cup(\cup_{\mathtt{k}=1}^{\mathtt{j}-1}V(P_{\mathtt{k}}))\cup\mathcal{L}(S)$ for $1\leq\mathtt{j}\leq\mathtt{t}$.
  3. (3)
    $\big{(}V(P_{\mathtt{j}})\setminus\operatorname{EndP}(P_{\mathtt{j}})\big{)}\cap\big{(}V(H)\cup(\cup_{\mathtt{i}=1}^{\mathtt{m}}V(C_{\mathtt{i}}))\cup(\cup_{\mathtt{k}=1}^{\mathtt{j}-1}V(P_{\mathtt{k}}))\cup\mathcal{L}(S)\big{)}=\emptyset$ for $\mathtt{1}\leq\mathtt{j}\leq\mathtt{t}$.
  4. (4)
    $\mathtt{t}=|\mathcal{L}(S)\setminus V(H)|+\tau(S)-\tau(H)$.
Lemma A.3 ([CDGL24+], Corollary A.4).
For $H\subset S$, we can decompose $E(S)\setminus E(H)$ into $\mathtt{m}$ cycles ${C}_{\mathtt{1}},\ldots,{C}_{\mathtt{m}}$ and $\mathtt{t}$ paths ${P}_{\mathtt{1}},\ldots,{P}_{\mathtt{t}}$ for some $\mathtt{m},\mathtt{t}\geq 0$ such that the following hold:
  1. (1)
    ${C}_{\mathtt{1}},\ldots,{C}_{\mathtt{m}}$ are independent cycles in $S$.
  2. (2)
    $V(P_{\mathtt{j}})\cap\big{(}V(H)\cup(\cup_{\mathtt{i}=1}^{\mathtt{m}}V(C_{\mathtt{i}}))\cup(\cup_{\mathtt{k}\neq\mathtt{j}}V(P_{\mathtt{k}}))\cup\mathcal{L}(S)\big{)}=\operatorname{EndP}(P_{\mathtt{j}})$ for $1\leq\mathtt{j}\leq\mathtt{t}$.
  3. (3)
    $\mathtt{t}\leq 5(|\mathcal{L}(S)\setminus V(H)|+\tau(S)-\tau(H))$.
Lemma A.4 ([CDGL24+], Lemma A.9).
For $H\subset\mathcal{K}_{n}$, we have (below we write $\mathfrak{P}=\{(p_{N+1},\ldots,p_{D}):\sum_{i=N+1}^{D}p_{i}\leq p,\text{ for all }N+1\leq l\leq D\}$ and $\sum_{\mathfrak{P}}$ for the summation over $(p_{N+1},\ldots,p_{D})\in\mathfrak{P}$)
#$\#\Big{\{}S\mbox{ admissible}\!:H\ltimes S;|\mathfrak{C}_{l}(S,H)|=0\mbox{ for }l>N;|\mathcal{L}(S)\setminus V(H)|+\tau(S)-\tau(H)=m;$
$|E(S)|-|E(H)|=p,|V(S)|-|V(H)|=q,|E(S)|\leq D\Big{\}}\leq(2D)^{3m}n^{q}\sum_{\mathfrak{P}}\prod_{j=N+1}^{D}\frac{1}{p_{j}!}\,.$ (A.1)
Lemma A.5 ([CDGL24+], Lemma A.10).
For $S\subset\mathcal{K}_{n}$ with $|E(S)|\leq D$, we have
[…]
$\mathfrak{C}_{j}(S;H)=m_{j},N+1\leq j\leq D\Big{\}}\leq D^{15m}\prod_{j=N+1}^{D}\binom{|\mathcal{C}_{j}(S)|}{m_{j}}\,.$
We now show some preliminary results on calculating expectations under $\sigma\sim\nu$.
Lemma A.6.
For a path $\mathcal{P}$ with $V(\mathcal{P})=\{v_{0},\ldots,v_{l}\}$ and $\operatorname{EndP}(\mathcal{P})=\{v_{0},v_{l}\}$, we have
$\mathbb{E}_{\sigma\sim\nu}\Big{[}\prod_{i=1}^{l}\Big{(}a+b\omega(\sigma_{i-1},\sigma_{i})\Big{)}\mid\sigma_{0},\sigma_{l}\Big{]}=a^{l}+b^{l}\cdot\omega(\sigma_{0},\sigma_{l})\,.$ (A.3)
Proof.
By independence, we see that $\mathbb{E}_{\sigma\sim\nu}\big{[}\prod_{i\in I}\omega(\sigma_{i-1},\sigma_{i})\mid\sigma_{0},\sigma_{l}\big{]}=0$ if $I\subsetneq[l]$. Thus,
$\mathbb{E}_{\sigma\sim\nu}\Big{[}\prod_{i=1}^{l}\Big{(}a+b\omega(\sigma_{i-1},\sigma_{i})\Big{)}\mid\sigma_{0},\sigma_{l}\Big{]}=a^{l}+b^{l}\mathbb{E}_{\sigma\sim\nu}\Big{[}\prod_{i=1}^{l}\omega(\sigma_{i-1},\sigma_{i})\mid\sigma_{0},\sigma_{l}\Big{]}\,.$
It remains to prove that
$\mathbb{E}_{\sigma\sim\nu}\Big{[}\prod_{i=1}^{l}\omega(\sigma_{i-1},\sigma_{i})\mid\sigma_{0},\sigma_{l}\Big{]}=\omega(\sigma_{0},\sigma_{l})\,.$ (A.4)
We shall show (A.4) by induction. The case $l=1$ follows immediately. Now we assume that (A.4) holds for $l$. Then we have
$\mathbb{E}_{\sigma\sim\nu}\Big{[}\prod_{i=1}^{l+1}\omega(\sigma_{i-1},\sigma_{i})\mid\sigma_{0},\sigma_{l+1}\Big{]}$
$=$ $\mathbb{E}_{\sigma\sim\nu}\Big{[}\omega(\sigma_{l},\sigma_{l+1})\mathbb{E}_{\sigma\sim\nu}\Big{[}\prod_{i=1}^{l}\omega(\sigma_{i-1},\sigma_{i})\mid\sigma_{0},\sigma_{l},\sigma_{l+1}\Big{]}\mid\sigma_{0},\sigma_{l+1}\Big{]}$
$=$ $\mathbb{E}_{\sigma\sim\nu}\Big{[}\omega(\sigma_{l},\sigma_{l+1})\omega(\sigma_{0},\sigma_{l})\mid\sigma_{0},\sigma_{l+1}\Big{]}=\omega(\sigma_{0},\sigma_{l+1})\,,$
which completes the induction procedure. ∎
Lemma A.7.
For $H\subset S$ with $\mathcal{L}(S)\not\subset V(H)$, we have
${}\mathbb{E}_{\sigma\sim\nu}\Big{[}\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\mid\{\sigma_{u}:u\in V(H)\}\Big{]}=0\,.$ (A.5)
Proof.
Denote $v\in\mathcal{L}(S)\setminus V(H)$ and $(v,w)\in E(S)$. Define $\sigma_{U}$ and $\sigma_{\setminus U}$ to be the restriction of $\sigma$ on $U$ and on $[n]\setminus U$, respectively. Also define $\nu_{U}$ and $\nu_{\setminus U}$ to be the restriction of $\nu$ on $U$ and on $[n]\setminus U$, respectively. Then we have (let $\mathtt{V}=V(H)$)
$\mathbb{E}_{\sigma\sim\nu}\Big{[}\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\mid\{\sigma_{u}:u\in V(H)\}\Big{]}$
=\ = $\mathbb{E}_{\sigma_{([n]\setminus\mathtt{V})}\sim\nu_{([n]\setminus\mathtt{V})}}\Big{[}\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\Big{]}$
=\ = $\mathbb{E}_{\sigma_{([n]\setminus(\mathtt{V}\cup\{v\}))}\sim\nu_{([n]\setminus(\mathtt{V}\cup\{v\}))}}\mathbb{E}_{\sigma_{\{v\}}\sim\nu_{\{v\}}}\Big{[}\omega(\sigma_{v},\sigma_{w})\prod_{(i,j)\in E(S)\setminus(E(H)\cup\{v,w\})}\omega(\sigma_{i},\sigma_{j})\Big{]}=0\,.\qed$
Appendix B Supplementary proofs
B.1 Proof of Theorem 2.4
The goal of this section is devoted to the proof of Theorem 2.4. We first briefly explain our proof ideas. Suppose on the contrary that there is an efficient algorithm $\mathcal{A}$ that takes $\mathsf{Y}$ as input and outputs either 00 or $1$ with
[…]
Since $\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime}),\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime})=o(1)$, we also have
${}\mathbb{P}^{\prime}(\mathcal{A}(\mathsf{Y})=1)=\Omega(1)\mbox{ and }\mathbb{Q}^{\prime}(\mathcal{A}(\mathsf{Y})=0)=1-\epsilon\mbox{ for some }\epsilon=\epsilon_{n}\to 0\,.$ (B.2)
Define
${}M=M_{n}=\max\big{\{}n,\lceil\epsilon_{n}^{-\frac{1}{2}}\rceil\big{\}}\to\infty\,.$ (B.3)
The crux of our argument is to consider the following hidden informative sample problem.
Definition B.1.
Consider the following hypothesis testing problem: we need to determine whether a sample $(\mathsf{Y}_{1},\mathsf{Y}_{2},\ldots,\mathsf{Y}_{M})$ where each $\mathsf{Y}_{i}\in\mathbb{R}^{N}$ is generated by
  • ¯$\overline{\mathcal{H}}_{0}$: we let $\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M}$ to be independently sampled from $\mathbb{Q}^{\prime}$.
  • ¯$\overline{\mathcal{H}}_{1}$: we first sample $\kappa\in\{1,\ldots,M\}$ uniformly at random, and (conditioned on the value of $\kappa$) we let $\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M}$ are independent samples with $\mathsf{Y}_{\kappa}$ generated from $\mathbb{P}^{\prime}$ and $\{\mathsf{Y}_{j}:j\neq\kappa\}$ generated from $\mathbb{Q}^{\prime}$.
In addition, denote $\overline{\mathbb{P}}$ and $\overline{\mathbb{Q}}$ to be the law of $(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M})$ under ¯$\overline{\mathcal{H}}_{1}$ and ¯$\overline{\mathcal{H}}_{0}$, respectively.
Now assuming that (B.2) holds, we see that
[…]
¯$\overline{\mathbb{P}}\Big{(}\big{(}\mathcal{A}(\mathsf{Y}_{1}),\ldots,\mathcal{A}(\mathsf{Y}_{M})\big{)}\neq(0,\ldots,0)\Big{)}\geq\Omega(1)\,.$ (B.5)
Thus, there is an efficient algorithm that achieves weak detection between $\overline{\mathbb{P}}$ and $\overline{\mathbb{Q}}$. Our next result, however, shows the low-degree advantage $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\overline{\mathbb{Q}}}\big{)}$ is bounded by $1+o(1)$.
Lemma B.2.
If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$, then $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\overline{\mathbb{Q}}}\big{)}=1+o(1)$.
Proof.
Note that
$\frac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\overline{\mathbb{Q}}}(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M})=\frac{1}{M}\sum_{i=1}^{M}\frac{\mathrm{d}\overline{\mathbb{P}}(\cdot\mid\kappa=i)}{\mathrm{d}\overline{\mathbb{Q}}}(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M})=\frac{1}{M}\sum_{i=1}^{M}\frac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}(\mathsf{Y}_{i})\,.$ (B.6)
In addition, recall (2.1). Define
$\mathcal{N}_{D}=\big{\{}f\in\mathcal{P}_{D}:f=0\mbox{ a.s. under }\mathbb{Q}^{\prime}\big{\}}\,.$
It is straightforward to verify that $\mathcal{N}_{D}$ is a linear subspace of $\mathcal{P}_{D}$. Also, consider the inner product $\langle f,g\rangle=\mathbb{E}_{\mathbb{Q}^{\prime}}[fg]$, denoting $\mathcal{N}_{D}^{\perp}$ to be the orthogonal space of $\mathcal{N}_{D}$ in $\mathcal{P}_{D}$, then $\mathcal{N}_{D}^{\perp}$ can be identified as a (finite-dimensional) Hilbert space where this inner product is non-degenerate (note that $\langle f,f\rangle>0$ for all $f\in\mathcal{N}_{D}^{\perp}\setminus\{0\}$). In addition, it is straightforward to check that $1\in\mathcal{N}_{D}^{\perp}$ with $\langle 1,1\rangle=1$. Thus, by applying the Schmidt orthogonalization procedure started with $1$, we can find a standard orthogonal basis $\{f_{\alpha}:\alpha\in\Lambda\}$ of the space $\mathcal{N}_{D}^{\perp}$ and with $0\in\Lambda$ and $f_{0}=1$. Note that in some cases (e.g., when $\mathbb{Q}^{\prime}$ is a product measure), this standard orthogonal basis have a closed form; however, for general $\mathbb{Q}^{\prime}$ the explicit form of $\{f_{\alpha}\}$ is often intractable. We first show that
${}\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=\Bigg{(}\sum_{\alpha\in\Lambda}\mathbb{E}_{\mathbb{P}^{\prime}}[f_{\alpha}(\mathsf{Y})]^{2}\Bigg{)}^{\frac{1}{2}}\,.$ (B.7)
Indeed, as $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$, we see that for all $f\in\mathcal{N}_{D}$ it must holds that $\mathbb{E}_{\mathbb{P}^{\prime}}[f]=0$ (since otherwise we will have $\frac{\mathbb{E}_{\mathbb{P}^{\prime}}[f]}{\mathbb{E}_{\mathbb{Q}^{\prime}}[f^{2}]}=\infty$). Thus, we have
$\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=\max_{f\in\mathcal{N}_{D}^{\perp}}\frac{\mathbb{E}_{\mathbb{P}^{\prime}}[f]}{\mathbb{E}_{\mathbb{Q}^{\prime}}[f^{2}]}\,.$
For any $f\in\mathcal{N}_{D}^{\perp}$, it can be uniquely expressed as
$f=\sum_{\alpha\in\Lambda}C_{\alpha}f_{\alpha}\,,$
where $C_{\alpha}$’s are real constants. Applying Cauchy-Schwartz inequality one gets
$\frac{\mathbb{E}_{\mathbb{P}^{\prime}}[f]}{\sqrt{\mathbb{E}_{\mathbb{Q}^{\prime}}[f^{2}]}}=\frac{\sum_{\alpha\in\Lambda}C_{\alpha}\mathbb{E}_{\mathbb{P}^{\prime}}[f_{\alpha}(\mathsf{Y})]}{\sqrt{\sum_{\alpha\in\Lambda}C_{\alpha}^{2}}}\leq\Bigg{(}\sum_{\alpha\in\Lambda}\mathbb{E}_{\mathbb{P}^{\prime}}[f_{\alpha}(\mathsf{Y})]^{2}\Bigg{)}^{1/2}\,,$
with equality holds if and only if $C_{\alpha}\propto\mathbb{E}_{\mathbb{P}^{\prime}}[f_{\alpha}]$. This yields (B.7). Now, note that ¯$\overline{\mathbb{Q}}=(\mathbb{Q}^{\prime})^{\otimes M}$ is a product measure of $\mathbb{Q}^{\prime}$, there is a natural standard orthogonal basis under $\overline{\mathbb{Q}}$, given by
$\Bigg{\{}\prod_{i=1}^{M}f_{\alpha_{i}}(\mathsf{Y}_{i}):\alpha_{i}\in\Lambda,\sum_{i=1}^{M}\operatorname{deg}(f_{\alpha_{i}})\leq D\Bigg{\}}\,.$
Thus, similarly as in (B.7), we see that
$\Big{(}\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\overline{\mathbb{Q}}}\big{)}\Big{)}^{2}=\sum_{\begin{subarray}{c}(\alpha_{1},\ldots,\alpha_{M}):\alpha_{i}\in\Lambda\\ \sum_{i=1}^{M}\operatorname{deg}(f_{\alpha_{i}})\leq D\end{subarray}}\mathbb{E}_{\overline{\mathbb{P}}}\Big{[}\prod_{i=1}^{M}f_{\alpha_{i}}(\mathsf{Y}_{i})\Big{]}^{2}\,.$ (B.8)
In addition, using (B.6), we see from direct calculation that
$\mathbb{E}_{\overline{\mathbb{P}}}\Big{[}\prod_{i=1}^{M}f_{\alpha_{i}}(\mathsf{Y}_{i})\Big{]}$ $=\mathbb{E}_{\overline{\mathbb{Q}}}\Big{[}\prod_{i=1}^{M}f_{\alpha_{i}}(\mathsf{Y}_{i})\cdot\frac{\mathrm{d}\overline{\mathbb{P}}}{\mathrm{d}\overline{\mathbb{Q}}}(\mathsf{Y}_{1},\ldots,\mathsf{Y}_{M})\Big{]}$
=(B.6)${=}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}_{\overline{\mathbb{Q}}}\Big{[}\prod_{i=1}^{M}f_{\alpha_{i}}(\mathsf{Y}_{i})\cdot\frac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}(\mathsf{Y}_{i})\Big{]}$
=$=\begin{cases}1\,,&(\alpha_{1},\ldots,\alpha_{M})=(0,\ldots,0)\,;\\ \frac{1}{M}\mathbb{E}_{\mathbb{P}^{\prime}}\big{[}f_{\alpha_{j}}(\mathsf{Y}_{j})\big{]}\,,&(\alpha_{1},\ldots,\alpha_{M})=(0,\ldots,0,\alpha_{j},0,\ldots,0)\,;\\ 0\,,&\mbox{otherwise}\,.\end{cases}$ (B.9)
Plugging (B.9) into (B.8), we get that
[…]
where in the second equality we use (B.7) and the assumption that $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$. This completes our proof. ∎
We can now finish the proof of Theorem 2.4.
Proof of Theorem 2.4.
Suppose on the contrary that there is an efficient algorithm $\mathcal{A}$ satisfying (B.1). Consider the hypothesis testing problem stated in Definition B.1. Using (B.4) and (B.5), we see that there is an efficient algorithm that achieves weak detection between $\overline{\mathbb{P}}$ and $\overline{\mathbb{Q}}$, which contradicts with Lemma B.2 and the equivalence between Items (1) and (2) in Conjecture 2.2. ∎
B.2 Proof of Lemma 3.5
In this subsection we provide the postponed proof of Lemma 3.5. For two graphs $S_{1},S_{2}\Subset\mathcal{K}_{n}$, define the polynomial $\phi_{S_{1},S_{2}}$ associated with $S_{1},S_{2}$ by
$\phi_{S_{1},S_{2}}\big{(}\{A_{i,j}\},\{B_{i,j}\}\big{)}=\big{(}q(1-q)\big{)}^{-\frac{|E(S_{1})|+|E(S_{2})|}{2}}\prod_{(i,j)\in E(S_{1})}\overline{A}_{i,j}\prod_{(i,j)\in E(S_{2})}\overline{B}_{i,j}\,,$ (B.10)
where $\overline{A}_{i,j}=A_{i,j}-q,\overline{B}_{i,j}=B_{i,j}-q$ for all $(i,j)\in\operatorname{U}$. In particular, $\phi_{\emptyset,\emptyset}\equiv 1$. It can be easily checked that $\mathcal{O}_{D}=\{\phi_{S_{1},S_{2}}:|E(S_{1})|+|E(S_{2})|\leq D\}$ constitutes a standard orthogonal basis of $\mathcal{P}_{D}$ under $\mathbb{Q}$. In addition, we say a polynomial $\phi_{S_{1},S_{2}}\in\mathcal{O}_{D}$ is admissible if both $S_{1}$ and $S_{2}$ are admissible graphs. Furthermore, we define $\mathcal{O}_{D}^{\prime}\subset\mathcal{O}_{D}$ as the set of admissible polynomials in $\mathcal{O}_{D}$, and define $\mathcal{P}_{D}^{\prime}\subset\mathcal{P}_{D}$ as the linear subspace spanned by polynomials in $\mathcal{O}_{D}^{\prime}$. It has been shown in [DDL23+, Proposition 3.4] that for any $f\in\mathcal{P}_{D}$, there exists $f^{\prime}\in\mathcal{P}_{D}^{\prime}$ such that $\mathbb{E}_{\mathbb{Q}}[(f^{\prime})^{2}]\leq 8\mathbb{E}_{\mathbb{Q}}[f^{2}]$ and $f^{\prime}=f$ a.s. under both $\overline{\mathbb{P}}$ and $\overline{\mathbb{P}}_{*}$. Thus, we get that
${}\mathsf{Adv}_{\leq D}\Big{(}\tfrac{\mathrm{d}\overline{\mathbb{P}}(\cdot\mid\pi_{*}(1)=i)}{\mathrm{d}\mathbb{Q}}\Big{)}=\sup_{f\in\mathcal{P}_{D}}\frac{\mathbb{E}_{\overline{\mathbb{P}}}[f\mid\pi_{*}(1)=1]}{\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}}\leq 2\sqrt{2}\sup_{f\in\mathcal{P}^{\prime}_{D}}\frac{\mathbb{E}_{\overline{\mathbb{P}}}[f\mid\pi_{*}(1)=1]}{\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}}\,.$ (B.11)
Thus, it suffices to show the right hand side of (B.11) is bounded by $O_{\delta}(1)$. Similar as (B.7), we have
$\sup_{f\in\mathcal{P}^{\prime}_{D}}\frac{\mathbb{E}_{\overline{\mathbb{P}}}[f\mid\pi_{*}(1)=1]^{2}}{\mathbb{E}_{\mathbb{Q}}[f^{2}]}=\sum_{\phi_{S_{1},S_{2}}\in\mathcal{O}^{\prime}_{D}}\mathbb{E}_{\overline{\mathbb{P}}}\big{[}\phi_{S_{1},S_{2}}\mid\pi_{*}(1)=i\big{]}^{2}\,.$
Without losing of generality, in the following we will only show that
${}\sum_{\phi_{S_{1},S_{2}}\in\mathcal{O}^{\prime}_{D}}\mathbb{E}_{\overline{\mathbb{P}}}\big{[}\phi_{S_{1},S_{2}}\mid\pi_{*}(1)=1\big{]}^{2}=O_{\delta}(1)\,.$ (B.12)
For a deterministic permutation $\pi\in\mathfrak{S}_{n}$, we use $\mathbb{P}_{\pi}$ and ¯$\overline{\mathbb{P}}_{\pi}$ to represent $\mathbb{P}_{*}(\cdot\mid\pi_{*}=\pi)$ and ¯$\overline{\mathbb{P}}_{*}(\cdot\mid\pi_{*}=\pi)$, respectively. For $S_{1},S_{2}\Subset\mathcal{K}_{n}$ with $|E(S_{1})|,|E(S_{2})|\leq D$, define
${}\mathtt{F}(S_{1},S_{2})=\sum_{\begin{subarray}{c}\mathbf{H}_{0}\in\mathcal{H}\\ \mathbf{H}_{0}\hookrightarrow S_{i},i=1,2\end{subarray}}n^{-\frac{|V(S_{1})|+|V(S_{2})|}{2}}\rho^{|E(\mathbf{H}_{0})|}d^{-6(|E(S_{1})|+|E(S_{2})|-2|E(\mathbf{H}_{0})|)}\operatorname{Aut}(\mathbf{H}_{0})\,,$ (B.13)
where the notation $\mathbf{H}\hookrightarrow S$ means that $\mathbf{H}$ can be embedded into $S$ as a subgraph. The key of our proof is the following estimation.
Proposition B.3.
We have
${}\big{|}\mathbb{E}_{\overline{\mathbb{P}}}[\phi_{S_{1},S_{2}}\mid\pi_{*}(1)=1]\big{|}\leq[1+o(1)]\big{(}\mathbf{1}_{1\not\in V(S_{1})\cap V(S_{2})}+n\mathbf{1}_{1\in V(S_{1})\cap V(S_{2})}\big{)}\cdot\mathtt{F}(S_{1},S_{2})\,.$ (B.14)
Proof.
For $S_{0}\hookrightarrow S_{1},S_{2}$, define
${}M(S_{0},S_{1},S_{2})=\rho^{|E(S_{0})|}n^{-\frac{|V(S_{1})|+|V(S_{2})|}{2}+|V(S_{0})|}D^{-7(|E(S_{1})|+|E(S_{2})|-2|E(S_{0})|)}\,.$ (B.15)
Using [DDL23+, Equation (3.29)], for any $S_{1},S_{2}\Subset\mathcal{K}_{n}$ with at most $D$ edges and any permutation $\pi\in\mathfrak{S}_{n}$, we have that (denote $S_{0}=S_{1}\cap S_{2}$ below)
${}\big{|}\mathbb{E}_{\overline{\mathbb{P}}_{\pi}}[\phi_{S_{1},S_{2}}]\big{|}\leq M(S_{0},S_{1},S_{2})\,.$ (B.16)
Thus, note that
$\big{|}\mathbb{E}_{\overline{\mathbb{P}}}[\phi_{S_{1},S_{2}}\mid\pi_{*}(1)=1]\big{|}$ $=\frac{1}{(n-1)!}\Big{|}\sum_{\pi\in\mathfrak{S}_{n},\pi(1)=1}\mathbb{E}_{\overline{\mathbb{P}}_{\pi}}[\phi_{S_{1},S_{2}}]\Big{|}$
$\leq\frac{1}{(n-1)!}\sum_{\pi\in\mathfrak{S}_{n},\pi(1)=1}\big{|}\mathbb{E}_{\overline{\mathbb{P}}_{\pi}}[\phi_{S_{1},S_{2}}]\big{|}\,,$ (B.17)
we can group the permutations $\pi\in\mathfrak{S}_{n}$ according to the realization of $S_{0}=S_{1}\cap\pi^{-1}(S_{2})$ and obtain that $\big{|}\mathbb{E}_{\overline{\mathbb{P}}}[\phi_{S_{1},S_{2}}\mid\pi_{*}(1)=1]\big{|}$ is bounded by
$\frac{1}{(n-1)!}\sum_{\begin{subarray}{c}S_{0},S^{\prime}_{0}:S_{0}\cong S_{0}^{\prime}\\ S_{0}\subset S_{1},S_{0}^{\prime}\subset S_{2}\end{subarray}}\#\{\pi\in\mathfrak{S}_{n}:\pi(1)=1,\pi(S_{0})=S_{0}^{\prime},S_{0}=S_{1}\cap\pi^{-1}(S_{2})\}M(S_{0},S_{1},S_{2})\,.$ (B.18)
For the case when $1\not\in V(S_{1})\cap V(S_{2})$, we must have $1\not\in V(S_{0})$ and thus
${}\mathsf{Enum}:=\#\{\pi\in\mathfrak{S}_{n}:\pi(1)=1,\pi(S_{0})=S_{0}^{\prime}\}$ (B.19)
is upper-bounded by
${}\operatorname{Aut}(S_{0})(n-1-|V(S_{0})|)!=[1+o(1)]\operatorname{Aut}(S_{0})(n-1)!\cdot n^{-|V(S_{0})|}\,.$ (B.20)
Thus, we have that (B.18) is further upper-bounded by (up to a factor of $(1+o(1))$)
$\sum_{\begin{subarray}{c}S_{0},S_{0}^{\prime}:S_{0}\cong S_{0}^{\prime}\\ S_{0}\subset S_{1},S_{0}^{\prime}\subset S_{2}\end{subarray}}n^{-|V(S_{0})|}\operatorname{Aut}(S_{0})M(S_{0},S_{1},S_{2})$
$=$ $\sum_{\begin{subarray}{c}\mathbf{H}_{0}\in\mathcal{H}\\ \mathbf{H}_{0}\xhookrightarrow{}S_{i},i=1,2\end{subarray}}\frac{\operatorname{Aut}(\mathbf{H}_{0})}{n^{|V(\mathbf{H}_{0})|}}M(\mathbf{H}_{0},S_{1},S_{2})\cdot\#\{(S_{0},S_{0}^{\prime}):S_{0}\subset S_{1},S_{0}^{\prime}\subset S_{2},S_{0}\cong S_{0}^{\prime}\cong\mathbf{H}_{0}\}$
$\leq$ $\sum_{\begin{subarray}{c}\mathbf{H}_{0}\in\mathcal{H}\\ \mathbf{H}_{0}\xhookrightarrow{}S_{i},i=1,2\end{subarray}}\frac{\operatorname{Aut}(\mathbf{H}_{0})}{n^{|V(\mathbf{H}_{0})|}}M(\mathbf{H}_{0},S_{1},S_{2})\cdot D^{|E(S_{1})|+E(S_{2})-2|E(\mathbf{H}_{0})|}\,,$
where the last inequality follows from Lemma A.1 (v) and the assumption that $|E(S_{1})|$ and $|E(S_{2})|$ are bounded by $D$. For the case $1\in V(S_{1})\cap V(S_{2})$, we have $\mathsf{Enum}\leq[1+o(1)]\operatorname{Aut}(S_{0})n!\cdot n^{-|V(S_{0})|}$, and thus we get an additional factor of $n$ in the final estimation. This completes the proof of Proposition B.3. ∎
We are now ready to establish (B.12), thus completing the proof of Lemma 3.5.
Proof of (B.12).
We divide the left hand side of (B.12) into two parts: the first part consists of those $S1,S2$ such that $1\not\in V(S_{1})\cap V(S_{2})$, and the second part consists of those with $1\in V(S_{1})\cap V(S_{2})$. We bound the first part
$\sum_{\begin{subarray}{c}\phi_{S_{1},S_{2}}\in\mathcal{O}^{\prime}_{D}\\ 1\not\in V(S_{1})\cap V(S_{2})\end{subarray}}\mathbb{E}_{\overline{\mathbb{P}}}\big{[}\phi_{S_{1},S_{2}}\mid\pi_{*}(1)=1\big{]}^{2}$ (B.21)
via Proposition B.3 by
$\ [1+o(1)]\sum_{\begin{subarray}{c}S_{1},S_{2}\Subset\mathcal{K}_{n}\text{admissible}\\ |E(S_{1})|+|E(S_{2})|\leq D\end{subarray}}\Bigg{(}\sum_{\begin{subarray}{c}\mathbf{H}_{0}\in\mathcal{H}\\ \mathbf{H}_{0}\xhookrightarrow{}S_{i},i=1,2\end{subarray}}\frac{\rho^{|E(\mathbf{H}_{0})|}\operatorname{Aut}(\mathbf{H}_{0})}{D^{6(|E(S_{1})|+|E(S_{2})|-2|E(\mathbf{H}_{0})|)}n^{\frac{1}{2}(|V(S_{1})|+|V(S_{2})|)}}\Bigg{)}^{2}$
$\leq$ $\ [1+o(1)]\sum_{\begin{subarray}{c}S_{1},S_{2}\Subset\mathcal{K}_{n}\text{admissible}\\ |E(S_{1})|+|E(S_{2})|\leq D\end{subarray}}\Bigg{(}\sum_{\begin{subarray}{c}\mathbf{H}_{0}\in\mathcal{H}\\ \mathbf{H}_{0}\xhookrightarrow{}S_{i},i=1,2\end{subarray}}\frac{\rho^{2|E(\mathbf{H}_{0})|}\operatorname{Aut}(\mathbf{H}_{0})^{2}}{D^{6(|E(S_{1})|+|E(S_{2})|-2|E(\mathbf{H}_{0})|)}n^{|V(S_{1})|+|V(S_{2})|}}\Bigg{)}$
×$\times\Bigg{(}\sum_{\begin{subarray}{c}\mathbf{H}_{0}\in\mathcal{H}\\ \mathbf{H}_{0}\xhookrightarrow{}S_{i},i=1,2\end{subarray}}D^{-6(|E(S_{1})|+|E(S_{2})|-2|E(\mathbf{H}_{0})|)}\Bigg{)}\,,$ (B.22)
where the last inequality comes form Cauchy-Schwartz inequality. It is straightforward to check that when $|E(S_{1})|+|E(S_{2})|\leq d$,
$\sum_{\begin{subarray}{c}\mathbf{H}_{0}\in\mathcal{H}\\ \mathbf{H}_{0}\xhookrightarrow{}S_{i},i=1,2\end{subarray}}d^{-6(|E(S_{1})|+|E(S_{2})|-2|E(\mathbf{H}_{0})|)}$
$\leq$ $\ \Big{(}\sum_{\begin{subarray}{c}\mathbf{H}_{0}\in\mathcal{H}\\ \mathbf{H}_{0}\xhookrightarrow{}S_{1}\end{subarray}}d^{-6(|E(S_{1})|-|E(\mathbf{H}_{0})|)}\Big{)}^{2}\leq\Big{(}\sum_{k=0}^{d}d^{-6k}\cdot d^{k}\Big{)}^{2}\leq 2\,,$ (B.23)
where the second inequality follows from Lemma A.1 (v). Plugging (B.23) into (B.21), we get that (B.21) is bounded by a constant factor times
$\sum_{\begin{subarray}{c}\mathbf{H}_{0}\in\mathcal{H}\\ \mathbf{H}_{0}\xhookrightarrow{}S_{i},i=1,2\end{subarray}}\rho^{2|E(\mathbf{H}_{0})|}\operatorname{Aut}(\mathbf{H}_{0})^{2}\sum_{\begin{subarray}{c}S_{1},S_{2}\Subset\mathcal{K}_{n}\text{admissible}\\ |E(S_{1})|+|E(S_{2})|\leq D\end{subarray}}\frac{n^{-|V(S_{1})|-|V(S_{2})|}}{D^{6(|E(S_{1})|+|E(S_{2})|-2|E(\mathbf{H}_{0})|)}}\,.$ (B.24)
By denoting $S_{1}\cong\mathbf{S}_{1}$ and $S_{2}\cong\mathbf{S}_{2}$ with $\mathbf{S}_{1},\mathbf{S}_{2}\in\mathcal{H}$, the right hand side of (B.24) reduces to
$\sum_{\mathbf{H}_{0}\in\mathcal{H}}\rho^{2|E(\mathbf{H}_{0})|}\operatorname{Aut}(\mathbf{H}_{0})^{2}\sum_{\begin{subarray}{c}\mathbf{S}_{1},\mathbf{S}_{2}\in\mathcal{H}\text{ admissible}\\ \mathbf{H}_{0}\xhookrightarrow{}\mathbf{S}_{i},i=1,2\\ |E(\mathbf{S}_{1})|+|E(\mathbf{S}_{2})|\leq d\end{subarray}}\frac{\#\{S_{1},S_{2}:S_{1}\cong\mathbf{S}_{1},S_{2}\cong\mathbf{S}_{2}\}}{D^{6(|E(\mathbf{S}_{1})|+|E(\mathbf{S}_{2})|-2|E(\mathbf{H}_{0})|)}n^{|V(\mathbf{S}_{1})|+|V(\mathbf{S}_{2})|}}$
$\leq$ $\sum_{\mathbf{H}_{0}\in\mathcal{H}\operatorname{admissible}}\rho^{2|E(\mathbf{H}_{0})|}\sum_{\begin{subarray}{c}\mathbf{S}_{1},\mathbf{S}_{2}\in\mathcal{H},\mathbf{H}_{0}\xhookrightarrow{}\mathbf{S}_{i},i=1,2\\ |E(\mathbf{S}_{1})|+|E(\mathbf{S}_{2})|\leq d\end{subarray}}\frac{\operatorname{Aut}(\mathbf{H}_{0})^{2}}{\operatorname{Aut}(\mathbf{S}_{1})\operatorname{Aut}(\mathbf{S}_{2})D^{6(|E(\mathbf{S}_{1})|+|E(\mathbf{S}_{2})|-2|E(\mathbf{H}_{0})|)}}\,,$ (B.25)
where the inequality follows from
#$\#\{S_{1},S_{2}:S_{1}\cong\mathbf{S}_{1},S_{2}\cong\mathbf{S}_{2}\}\leq[1+o(1)]\cdot\frac{n^{|V(\mathbf{S}_{1})|}}{\operatorname{Aut}(\mathbf{S}_{1})}\cdot\frac{n^{|V(\mathbf{S}_{2})|}}{\operatorname{Aut}(\mathbf{S}_{2})}\,.$
Using Lemma A.1 (iii), we have $\frac{\operatorname{Aut}(\mathbf{H}_{0})^{2}}{\operatorname{Aut}(\mathbf{S}_{1})\operatorname{Aut}(\mathbf{S}_{2})}\leq d^{2(|E(\mathbf{S}_{1})|+|E(\mathbf{S}_{2})|-2|E(\mathbf{H}_{0})|)}$, and thus the right hand side of (B.25) is upper-bounded by
$\ \sum_{\mathbf{H}_{0}\in\mathcal{H}\operatorname{admissible}}\rho^{2|E(\mathbf{H}_{0})|}\sum_{\begin{subarray}{c}\mathbf{S}_{1},\mathbf{S}_{2}\subset\mathcal{H},\mathbf{H}_{0}\xhookrightarrow{}\mathbf{S}_{i},i=1,2\\ |E(\mathbf{S}_{1})|+|E(\mathbf{S}_{2})|\leq d\end{subarray}}d^{-4(|E(\mathbf{S}_{1})|+|E(\mathbf{S}_{2})|-2|E(\mathbf{H}_{0})|)}$ (B.26)
$\leq$ $\ 4\sum_{\begin{subarray}{c}\mathbf{H}_{0}\in\mathcal{H}\operatorname{admissible}\\ |E(\mathbf{H}_{0})|\leq d\end{subarray}}\rho^{2|E(\mathbf{H}_{0})|}\,,$
where we used the fact that (similar to (B.23))
$\ \sum_{\begin{subarray}{c}\mathbf{S}_{1},\mathbf{S}_{2}\in\mathcal{H},\mathbf{H}_{0}\xhookrightarrow{}\mathbf{S}_{i},i=1,2\\ |E(\mathbf{S}_{1})|+|E(\mathbf{S}_{2})|\leq d\end{subarray}}d^{-4(|E(\mathbf{S}_{1})|+|E(\mathbf{S}_{2})|-2|E(\mathbf{H}_{0})|)}$
$\leq$ $\ \Bigg{(}\sum_{\begin{subarray}{c}\mathbf{H}_{0}\xhookrightarrow{}\mathbf{S},|E(\mathbf{S})|\leq d\end{subarray}}d^{-4(|E(\mathbf{S})|-|E(\mathbf{H})|)}\Bigg{)}^{2}\leq\Bigg{(}\sum_{k=0}^{d}d^{-4k}\cdot(2d)^{2k}\Bigg{)}^{2}\leq 4\,.$
In conclusion, we have shown that
(B.25)$\leq O_{\delta}(1)\cdot\sum_{\mathbf{H}_{0}\in\mathcal{H}\operatorname{admissible}}\rho^{2|E(\mathbf{H}_{0})|}\rho^{2|E(\mathbf{H}_{0})|}=O_{\delta}(1)\,,$ (B.27)
where the equality follows from [DDL23+, Lemma A.3]. Thus, we have
(B.21)$=O_{\delta}(1)\,.$ ) . (B.28)
We now deal with the second part
${}\sum_{\begin{subarray}{c}\phi_{S_{1},S_{2}}\in\mathcal{O}^{\prime}_{D}\\ 1\in V(S_{1})\cap V(S_{2})\end{subarray}}\mathbb{E}_{\overline{\mathbb{P}}}\big{[}\phi_{S_{1},S_{2}}\mid\pi_{*}(1)=1\big{]}^{2}\,.$ (B.29)
Following a similar derivation of (B.24), we see that (B.29) is bounded by a constant factor times
$\sum_{\begin{subarray}{c}\mathbf{H}_{0}\in\mathcal{H}\\ \mathbf{H}_{0}\xhookrightarrow{}S_{i},i=1,2\end{subarray}}\rho^{2|E(\mathbf{H}_{0})|}\operatorname{Aut}(\mathbf{H}_{0})^{2}\sum_{\begin{subarray}{c}S_{1},S_{2}\Subset\mathcal{K}_{n}\text{admissible}\\ 1\in V(S_{1})\cap V(S_{2})\\ |E(S_{1})|+|E(S_{2})|\leq D\end{subarray}}\frac{n^{-|V(S_{1})|-|V(S_{2})|}}{D^{6(|E(S_{1})|+|E(S_{2})|-2|E(\mathbf{H}_{0})|)}}\times n^{2}$
(Note that in the above there is an additional factor of $n^{2}$ compared with (B.24).) However, due to the restriction $1\in V(S_{1})\cap V(S_{2})$, for any fixed $\mathbf{S}_{1},\mathbf{S}_{2}\in\mathcal{H}$ the enumeration of $(S_{1},S_{2})$ such that $S_{1}\cong\mathbf{S}_{1},S_{2}\cong\mathbf{S}_{2}$ is upper-bounded by
$\frac{n^{|V(\mathbf{S}_{1})|-1}}{\operatorname{Aut}(\mathbf{S}_{1})}\cdot\frac{n^{|V(\mathbf{S}_{2})|-1}}{\operatorname{Aut}(\mathbf{S}_{2})}=\frac{n^{|V(\mathbf{S}_{1})|+|V(\mathbf{S}_{2})|}}{\operatorname{Aut}(\mathbf{S}_{1})\operatorname{Aut}(\mathbf{S}_{2})}\times n^{-2}\,.$
Thus, after canceling the factor of $n^{-2}$ with the factor of $n^{2}$ above we derive the same upper bound as we obtain on (B.21). This shows that
(B.29)$=O_{\delta}(1)\,.$ ) . (B.30)
Combining (B.28) and (B.30) leads to the desired result. ∎
B.3 Proof of Lemma 3.9
Now it remains to prove Lemma 3.9. Recall the definition of $\phi_{S_{1},S_{2}}$ in (B.10) (here we will take $q=\tfrac{\lambda s}{n}$). We first show the following estimation. Again, we say a polynomial $\phi_{S_{1},S_{2}}\in\mathcal{O}_{D}$ is admissible if both $S_{1}$ and $S_{2}$ are admissible graphs. Furthermore, we define $\mathcal{O}^{\prime}_{D}\subset\mathcal{O}_{D}$ as the set of admissible polynomials in $\mathcal{O}_{D}$, and define $\mathcal{P}^{\prime}_{D}\subset\mathcal{P}_{D}$ as the linear subspace spanned by polynomials in $\mathcal{O}^{\prime}_{D}$. Using [CDGL24+, Proposition 4.5], we see that for any $f\in\mathcal{P}_{D}$, there exists some $f^{\prime}\in\mathcal{P}_{D}^{\prime}$ such that $\mathbb{E}_{\mathbb{Q}}[(f^{\prime})^{2}]\leq O_{\delta,k}(1)\cdot\mathbb{E}_{\mathbb{Q}}[f^{2}]$ and $f^{\prime}=f$ a.s. under both $\widetilde{\mathbb{P}}_{*}$ and $\widetilde{\mathbb{P}}$. Thus, we get that
$\mathsf{Adv}_{\leq D}\Big{(}\tfrac{\mathrm{d}\widetilde{\mathbb{P}}(\cdot\mid\pi_{*}(i)=j)}{\mathrm{d}\mathbb{Q}}\Big{)}=\sup_{f\in\mathcal{P}_{D}}\frac{\mathbb{E}_{\widetilde{\mathbb{P}}}[f\mid\pi_{*}(i)=j]}{\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}}\leq O_{\delta,k}(1)\cdot\sup_{f\in\mathcal{P}^{\prime}_{D}}\frac{\mathbb{E}_{\widetilde{\mathbb{P}}}[f\mid\pi_{*}(i)=j]}{\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}}\,.$
Thus, it suffices to show that
$\sup_{f\in\mathcal{P}^{\prime}_{D}}\frac{\mathbb{E}_{\widetilde{\mathbb{P}}}[f\mid\pi_{*}(1)=1]}{\sqrt{\mathbb{E}_{\mathbb{Q}}[f^{2}]}}=\sum_{\phi_{S_{1},S_{2}}\in\mathcal{O}^{\prime}_{D}}\mathbb{E}_{\widetilde{\mathbb{P}}}\big{[}\phi_{S_{1},S_{2}}\mid\pi_{*}(i)=j\big{]}^{2}=O_{\delta,k}(1)\,.$
For notational simplicity, in the following we will only show
${}\sum_{\phi_{S_{1},S_{2}}\in\mathcal{O}^{\prime}_{D}}\mathbb{E}_{\widetilde{\mathbb{P}}}\big{[}\phi_{S_{1},S_{2}}\mid\pi_{*}(1)=1\big{]}^{2}=O_{\delta,k}(1)\,,$ (B.31)
and the general cases for $\pi_{*}(i)=j$ can be derived in a similar manner. For $H\subset S$, we define $\mathtt{N}(S,H)$ to be
$\mathtt{N}(S,H)$ $=\big{(}\tfrac{D^{28}}{n^{0.1}}\big{)}^{\frac{1}{2}(|\mathcal{L}(S)\setminus V(H)|+\tau(S)-\tau(H))}(1-\tfrac{\delta}{2})^{|E(S)|-|E(H)|}\,.$ (B.32)
We first show the following estimation.
Proposition B.4.
For all admissible $S_{1},S_{2}\Subset\mathcal{K}_{n}$ with $|E(S_{1})|,|E(S_{2})|\leq D$, we have that $\big{|}\mathbb{E}_{\widetilde{\mathbb{P}}}[\phi_{S_{1},S_{2}}\mid\pi_{*}(1)=1]\big{|}$ is bounded by $O_{\delta,k}(1)$ times (note that in the summation below $H_{1},H_{2}$ may have isolated vertices)
${}\sum_{\begin{subarray}{c}H_{1}\subset S_{1},H_{2}\subset S_{2}\\ H_{1}\cong H_{2}\end{subarray}}\frac{(\sqrt{\alpha}-\tfrac{\delta}{2})^{|E(H_{1})|}\operatorname{Aut}(H_{1})}{n^{|V(H_{1})|-\mathbf{1}_{\{1\in V(H_{1}\cap V(H_{2}))\}}}}*\frac{2^{|\mathfrak{C}(S_{1},H_{1})|+|\mathfrak{C}(S_{2},H_{2})|}\mathtt{N}(S_{1},H_{1})\mathtt{N}(S_{2},H_{2})}{n^{\frac{1}{2}(|V(S_{1})|+|V(S_{2})|-|V(H_{1})|-|V(H_{2})|)}}\,.$ (B.33)
Proof.
For each deterministic permutation $\pi\in\mathfrak{S}_{n}$ and each labeling $\sigma\in[k]^{n}$, we denote ~$\widetilde{\mathbb{P}}_{\sigma,\pi}=\widetilde{\mathbb{P}}(\cdot\mid\sigma_{*}=\sigma,\pi_{*}=\pi)$, ~$\widetilde{\mathbb{P}}_{\pi}=\widetilde{\mathbb{P}}(\cdot\mid\pi_{*}=\pi)$ and ~$\widetilde{\mathbb{P}}_{\sigma}=\widetilde{\mathbb{P}}(\cdot\mid\sigma_{*}=\sigma)$ respectively. It is clear that
$\Big{|}\mathbb{E}_{\widetilde{\mathbb{P}}}\big{[}\phi_{S_{1},S_{2}}\mid\pi_{*}(1)=1\big{]}\Big{|}$ $=\Big{|}\frac{1}{(n-1)!}\sum_{\begin{subarray}{c}\pi\in\mathfrak{S}_{n}\\ \pi(1)=1\end{subarray}}\mathbb{E}_{\widetilde{\mathbb{P}}_{\pi}}[\phi_{S_{1},S_{2}}]\Big{|}$
$\leq\frac{1}{(n-1)!}\sum_{\begin{subarray}{c}\pi\in\mathfrak{S}_{n}\\ \pi(1)=1\end{subarray}}\big{|}\mathbb{E}_{\widetilde{\mathbb{P}}_{\pi}}[\phi_{S_{1},S_{2}}]\big{|}\,.$ (B.34)
For $H\subset S$, we define
${}\mathtt{M}(S,H)=\big{(}\tfrac{D^{8}}{n^{0.1}}\big{)}^{\frac{1}{2}(|\mathcal{L}(S)\setminus V(H)|+\tau(S)-\tau(H))}(1-\tfrac{\delta}{2})^{|E(S)|-|E(H)|}\,.$ (B.35)
Applying [CDGL24+, Lemma 4.10], we have that for all admissible $S_{1},S_{2}\Subset\mathcal{K}_{n}$ with at most $D$ edges and for all permutation $\pi$ on $[n]$, denote $H_{1}=S_{1}\cap\pi^{-1}(S_{2})$ and $H_{2}=\pi(S_{1})\cap S_{2}$. We have that $\big{|}\mathbb{E}_{\widetilde{\mathbb{P}}_{\pi}}[\phi_{S_{1},S_{2}}]\big{|}$ is bounded by $O_{\delta,k}(1)$ times
$(\sqrt{\alpha}-\tfrac{\delta}{2})^{|E(H_{1})|}\sum_{H_{1}\ltimes K_{1}\subset S_{1}}\sum_{H_{2}\ltimes K_{2}\subset S_{2}}\frac{\mathtt{M}(S_{1},K_{1})\mathtt{M}(S_{2},K_{2})\mathtt{M}(K_{1},H_{1})\mathtt{M}(K_{2},H_{2})}{n^{\frac{1}{2}(|V(S_{1})|+|V(S_{2})|-|V(H_{1})|-|V(H_{2})|)}}\,.$ (B.36)
Note that we have
${}\mu\big{(}\{\pi:\pi(1)=1,\pi(H_{1})=H_{2}\}\big{)}\leq[1+o(1)]\cdot\operatorname{Aut}(H_{1})n^{-|V(H_{1})|-1+\mathbf{1}_{1\in V(H_{1})\cap V(H_{2})}}\,.$ (B.37)
Thus, it yields that the right-hand side of (B.34) is bounded by (up to a $O_{\delta,k}(1)$ factor)
$\sum_{\begin{subarray}{c}H_{1},H_{2}:H_{1}\cong H_{2}\\ H_{1}\subset S_{1},H_{2}\subset S_{2}\end{subarray}}\frac{\operatorname{Aut}(H_{1})(\sqrt{\alpha}-\tfrac{\delta}{2})^{|E(H_{1})|}}{n^{|V(H_{1})|-\mathbf{1}_{1\in V(H_{1})\cap V(H_{2})}}}*\frac{\mathtt{P}(S_{1},H_{1})\mathtt{P}(S_{2},H_{2})}{n^{\frac{1}{2}(|V(S_{1})|+|V(S_{2})|-|V(H_{1})|-|V(H_{2})|)}}\,,$ (B.38)
where (recall (B.35))
${}\mathtt{P}(S,H)=\sum_{H\ltimes K\subset S}\mathtt{M}(S,K)\mathtt{M}(K,H)\,.$ (B.39)
Using [CDGL24+, Claim 4.11], we have that $\mathtt{P}(S,H)\leq[1+o(1)]\cdot 2^{|\mathfrak{C}(S,H)|}\mathtt{N}(S,H)$, thus leading to the desired result. ∎
Now we prove (B.31) formally, thus finishing the proof of Lemma 3.9.
Proof of (B.31).
By Proposition B.4 and Cauchy-Schwartz inequality, (B.31) is upper-bounded by $O_{\delta,k}(1)$ times
$\sum_{\phi_{S_{1},S_{2}}\in\mathcal{O}^{\prime}_{D}}\Big{(}\sum_{\begin{subarray}{c}H_{1}\subset S_{1},H_{2}\subset S_{2}\\ H_{1}\cong H_{2}\end{subarray}}n^{0.02|\mathcal{I}(H_{1})|}\mathtt{N}(S_{1},H_{1})\mathtt{N}(S_{2},H_{2})\Big{)}\times$ (B.40)
$\Big{(}\sum_{\begin{subarray}{c}H_{1}\subset S_{1},H_{2}\subset S_{2}\\ H_{1}\cong H_{2}\end{subarray}}\frac{(\sqrt{\alpha}-\tfrac{\delta}{2})^{2|E(H_{1})|}\operatorname{Aut}(H_{1})^{2}}{n^{2|V(H_{1})|-2\cdot\mathbf{1}_{1\in V(H_{1})\cap V(H_{2})}+0.02|\mathcal{I}(H_{1})|}}\cdot\frac{4^{|\mathfrak{C}(S_{1},H_{1})|+|\mathfrak{C}(S_{2},H_{2})|}\mathtt{N}(S_{1},H_{1})\mathtt{N}(S_{2},H_{2})}{n^{(|V(S_{1})|+|V(S_{2})|-|V(H_{1})|-|V(H_{2})|)}}\Big{)}\,.$
It was shown in [CDGL24+, Equation (4.35)] that the first bracket in (B.40) is bounded by $O(1)$. Thus, we see that the left hand side of (B.31) is bounded by $O_{\delta,k}(1)$ times
$\sum_{\begin{subarray}{c}H_{1}\cong H_{2},H_{1},H_{2}\text{ admissible}\\ |E(H_{1})|+|E(H_{2})|\leq D\end{subarray}}$ $\frac{(\sqrt{\alpha}-\tfrac{\delta}{2})^{2|E(H_{1})|}\operatorname{Aut}(H_{1})^{2}}{n^{2|V(H_{1})|-2\cdot\mathbf{1}_{1\in V(H_{1})\cap V(H_{2})}+0.02|\mathcal{I}(H_{1})|}}\cdot\mathtt{L}(H_{1},H_{2})\,,$
where
$\mathtt{L}(H_{1},H_{2})=\sum_{\begin{subarray}{c}H_{1}\subset S_{1},H_{2}\subset S_{2}\end{subarray}}\frac{4^{|\mathfrak{C}(S_{1},H_{1})|+|\mathfrak{C}(S_{2},H_{2})|}\mathtt{N}(S_{1},H_{1})\mathtt{N}(S_{2},H_{2})}{n^{(|V(S_{1})|+|V(S_{2})|-|V(H_{1})|-|V(H_{2})|)}}\,.$
Using [CDGL24+, Equation (4.37)], we see that $\mathtt{L}(H_{1},H_{2})=O_{\delta,k}(1)$. Thus, the left hand side of (B.31) is bounded by $O_{\delta,k}(1)$ times
$\sum_{\begin{subarray}{c}H_{1}\cong H_{2},H_{1},H_{2}\text{ admissible}\\ |E(H_{1})|+|E(H_{2})|\leq D\end{subarray}}\frac{(\sqrt{\alpha}-\tfrac{\delta}{2})^{2|E(H_{1})|}\operatorname{Aut}(H_{1})^{2}}{n^{2|V(H_{1})|-2\cdot\mathbf{1}_{1\in V(H_{1})\cap V(H_{2})}+0.02|\mathcal{I}(H_{1})|}}$
=\ = $\sum_{\begin{subarray}{c}H_{1}\cong H_{2},H_{1},H_{2}\text{ admissible}\\ |E(H_{1})|+|E(H_{2})|\leq D\\ 1\not\in V(H_{1})\cap V(H_{2})\end{subarray}}\frac{(\sqrt{\alpha}-\tfrac{\delta}{2})^{2|E(H_{1})|}\operatorname{Aut}(H_{1})^{2}}{n^{2|V(H_{1})|+0.02|\mathcal{I}(H_{1})|}}$ (B.41)
$+$ $\sum_{\begin{subarray}{c}H_{1}\cong H_{2},H_{1},H_{2}\text{ admissible}\\ |E(H_{1})|+|E(H_{2})|\leq D\\ 1\in V(H_{1})\cap V(H_{2})\end{subarray}}\frac{(\sqrt{\alpha}-\tfrac{\delta}{2})^{2|E(H_{1})|}\operatorname{Aut}(H_{1})^{2}}{n^{2|V(H_{1})|+0.02|\mathcal{I}(H_{1})|}}\times n^{2}\,.$ (B.42)
Recall that we use $\widetilde{H}_{1}$ to denote the subgraph of $H_{1}$ obtained by removing all the vertices in $\mathcal{I}(H_{1})$. In addition, for $|V(H_{1})|\leq|V(S_{1})|\leq 2D$, we have $\operatorname{Aut}(H_{1})=\operatorname{Aut}(\widetilde{H}_{1})\cdot|\mathcal{I}(H_{1})|!\leq(2D)^{|\mathcal{I}(H_{1})|}\operatorname{Aut}(\widetilde{H}_{1})$. Thus, we see that Thus, we have that
[…]
$\circeq\sum_{\begin{subarray}{c}|E(\mathbf{H})|\leq D,\mathcal{I}(\mathbf{H})=\emptyset\\ \mathbf{H}\textup{ is admissible}\end{subarray}}(\sqrt{\alpha}-\tfrac{\delta}{4})^{2|E(\mathbf{H})|}\leq O_{\delta,k}(1)\,,$
As for (B.42), due to the restriction $1\in V(H_{1})\cap V(H_{2})$, for any fixed $\mathbf{H}_{1},\mathbf{H}_{2}\in\mathcal{H}$ the enumeration of $(H_{1},H_{2})$ such that $H_{1}\cong\mathbf{H}_{1},H_{2}\cong\mathbf{H}_{2}$ is upper-bounded by
$\frac{n^{|V(\mathbf{H}_{1})|-1}}{\operatorname{Aut}(\mathbf{H}_{1})}\cdot\frac{n^{|V(\mathbf{H}_{2})|-1}}{\operatorname{Aut}(\mathbf{H}_{2})}=\frac{n^{|V(\mathbf{H}_{1})|+|V(\mathbf{H}_{2})|}}{\operatorname{Aut}(\mathbf{H}_{1})\operatorname{Aut}(\mathbf{H}_{2})}\times n^{-2}\,.$
Thus, after canceling the factor of $n^{-2}$ with the factor of $n^{2}$ above we derive the same upper bound as we obtain on (B.42). This yields that (B.42)$=O_{\delta,k}(1)$ ) and thus completes our proof. ∎
B.4 Proof of Lemma 4.4
In this subsection we present the proof of Lemma 4.4. The crucial input in our proof of Lemma 4.4 is the following estimation on $\Xi(H)$.
Lemma B.5.
Assuming (4.3) with $|E(S)|\leq D$ and $\mathcal{L}(S)=\emptyset$. We have the following estimation:
  1. (1)
    If $S=S_{1}\cup S_{2}$ such that $V(S_{1})\cap V(S_{2})=\emptyset$, then $\Xi(S)=\Xi(S_{1})\Xi(S_{2})$.
  2. (2)
    If $S$ can be decomposed into different cycles $C_{1},\ldots,C_{m}$, then
    $|\Xi(S)|\leq k^{m}(1-\delta/2)^{\frac{|E(H)|}{2}}n^{-\frac{|E(H)|}{2}}\,.$
  3. (3)
    If $\tau(S)>0$ and $\mathfrak{C}_{k}(S)=m_{k}$ for $3\leq k\leq D$, then
    $|\Xi(S)|\leq(10\tau(S))!\cdot(2kD)^{10\tau(H)}(1-\delta/2)^{\frac{|E(H)|}{2}}n^{-\frac{|E(H)|}{2}}\,.$
Proof.
We first prove Item (1). We will prove by induction on $|E(S)|$. The case $S=\emptyset$ is trivial. Now Suppose $S=S_{1}\cup S_{2}$, for all $H\subset S$ with $\mathcal{L}(H)=\emptyset$, we have $H=H_{1}\cup H_{2}$ with $H_{1}\subset S_{1},H_{2}\subset S_{2}$ and $\mathcal{L}(H_{1}),\mathcal{L}(H_{2})=\emptyset$. In addition, denote
$\mathtt{P}(S)=\mathbb{E}_{\sigma\sim\nu}\Big{[}\prod_{(i,j)\in E(S)}\mathtt{h}(\sigma_{i},\sigma_{j})\Big{]}\,,$
$\mathtt{Q}(S,H)=\mathbb{E}_{\sigma\sim\nu}\Big{[}\prod_{(i,j)\in E(H)}\mathtt{h}(\sigma_{i},\sigma_{j})\prod_{(i,j)\in E(S)\setminus E(H)}\omega(\sigma_{i},\sigma_{j})\Big{]}\,.$
We have $\mathtt{P}(S)=\mathtt{P}(S_{1})\mathtt{P}(S_{2})$ and $\mathtt{Q}(S,H)=\mathtt{Q}(S_{1},H_{1})\mathtt{Q}(S_{2},H_{2})$. Also, it is straightforward to verify that $\mathtt{Q}(S,S)=\mathtt{P}(S)$. Thus, we have
$-\Xi(S)=$ $\sum_{\begin{subarray}{c}H_{1}\subsetneq S_{1},H_{2}\subsetneq S_{2}\\ \mathcal{L}(H_{1}),\mathcal{L}(H_{2})=\emptyset\end{subarray}}\frac{\mathtt{Q}(S_{1},H_{1})\mathtt{Q}(S_{2},H_{2})\Xi(H_{1}\sqcup H_{2})}{\mathtt{P}(S_{1})\mathtt{P}(S_{2})}+\sum_{\begin{subarray}{c}H_{2}\subsetneq S_{2}\\ \mathcal{L}(H_{2})=\emptyset\end{subarray}}\frac{\mathtt{Q}(S_{2},H_{2})\Xi(S_{1}\sqcup H_{2})}{\mathtt{P}(S_{2})}$
$+\sum_{\begin{subarray}{c}H_{1}\subsetneq S_{1}\\ \mathcal{L}(H_{1})=\emptyset\end{subarray}}\frac{\mathtt{Q}(S_{1},H_{1})\Xi(H_{1}\sqcup S_{2})}{\mathtt{P}(S_{1})}$
$=$ $\sum_{\begin{subarray}{c}H_{1}\subsetneq S_{1},H_{2}\subsetneq S_{2}\\ \mathcal{L}(H_{1}),\mathcal{L}(H_{2})=\emptyset\end{subarray}}\frac{\mathtt{Q}(S_{1},H_{1})\mathtt{Q}(S_{2},H_{2})\Xi(H_{1})\Xi(H_{2})}{\mathtt{P}(S_{1})\mathtt{P}(S_{2})}+\sum_{\begin{subarray}{c}H_{2}\subsetneq S_{2}\\ \mathcal{L}(H_{2})=\emptyset\end{subarray}}\frac{\mathtt{Q}(S_{2},H_{2})\Xi(S_{1})\Xi(H_{2})}{\mathtt{P}(S_{2})}$
$+\sum_{\begin{subarray}{c}H_{1}\subsetneq S_{1}\\ \mathcal{L}(H_{1})=\emptyset\end{subarray}}\frac{\mathtt{Q}(S_{1},H_{1})\Xi(H_{1})\Xi(S_{2})}{\mathtt{P}(S_{1})}$
$=$ $\ (-\Xi(S_{1}))\cdot(-\Xi(S_{2}))+\Xi(S_{1})\cdot(-\Xi(S_{2}))+(-\Xi(S_{1}))\cdot\Xi(S_{2})=-\Xi(S_{1})\Xi(S_{2})\,,$
where the second equality follows from induction hypothesis and the third equality follows from (4.22). This yields Item (1).
Now we focus on Item (2). Based on Item (1). it suffices to show that for any cycle $C$ with $|E(C)|\leq D$ we have
${}|\Xi(C)|\leq k(1-\delta)^{\frac{|E(C)|}{2}}n^{-\frac{|E(C)|}{2}}\,.$ (B.43)
Note that using (4.21) and (4.22), we have that
$\Xi(C)=\tfrac{\mathtt{Q}(S,\emptyset)}{\mathtt{P}(S)}\cdot\big{(}\tfrac{\epsilon^{2}\lambda}{n}\big{)}^{|E(C)|/2}=\tfrac{(k-1)}{\mathtt{P}(S)}\cdot(\epsilon^{2}\lambda)^{|E(C)|/2}n^{-|E(C)|/2}\,,$
where in the second equality we use Lemma A.6. Note that since $\langle\sigma_{i},\sigma_{j}\rangle\in\{-1,k-1\}$, we have
$\mathtt{h}(\sigma_{i},\sigma_{j})=a\omega(\sigma_{i},\sigma_{j})+b\,,$
where
$a=[1+o(1)]\cdot\tfrac{(k-1)\sqrt{1-\epsilon}+\sqrt{1+\epsilon(k-1)}}{k}\mbox{ and }b=[1+o(1)]\cdot\tfrac{\sqrt{1+\epsilon(k-1)}-\sqrt{1-\epsilon}}{k}\,.$
Thus, using Lemma A.6 we get that
$\mathtt{P}(C)=a^{|E(C)|}+(k-1)b^{|E(C)|}\geq a^{|E(C)|}\,.$
Plugging this result into the bound on $\Xi(C)$ yields Item (2) (recall (4.2)).
Finally we prove Item (3). Again based on Item (1), we may assume that $S$ is connected. We prove Item (3) by induction on $\tau(S)$. The case where $\tau(S)=0$ is proved in Item (2). Now assume Item (3) holds for $\tau(S)\leq m-1$, we now focus on $\tau(S)=m$. Using Lemma A.3, we can decompose $S$ into $\mathtt{t}$ disjoint self-avoiding paths $P_{\mathtt{1}},\ldots,P_{\mathtt{t}}$ with $\mathtt{t}\leq 3m$ satisfying Lemma A.3. In addition, applying Lemma A.3 for $H\subset S$ and $\mathcal{L}(H)=\mathcal{L}(S)=\emptyset$, we see that assuming $\tau(H)=m-r$ for some $r>0$, we then have $S\setminus H=P_{\mathtt{i}_{1}}\cup\ldots\cup\mathtt{P}_{\mathtt{i}_{s}}$ where $s\leq 3r$. Denote $\mathtt{V}=\cup_{1\leq\mathtt{i}\leq m}\operatorname{EndP}(P_{\mathtt{i}})$, by conditioning on $\{\sigma_{u}:u\in\mathtt{V}\}$ we see that
$\Big{\{}\prod_{(i,j)\in E(P_{\mathtt{i}})}\mathtt{h}(\sigma_{i},\sigma_{j}):1\leq\mathtt{i}\leq m\Big{\}}$
are conditional independent, with (denote $\operatorname{EndP}(P_{\mathtt{i}})=\{u_{\mathtt{i}},v_{\mathtt{i}}\}$)
$\mathbb{E}\Big{[}\prod_{(i,j)\in E(P_{\mathtt{i}})}\mathtt{h}(\sigma_{i},\sigma_{j})\mid\{\sigma_{u}:u\in\mathtt{V}\}\Big{]}=a^{|E(P_{\mathtt{i}})|}+b^{|E(P_{\mathtt{i}})|}\omega(u_{\mathtt{i}},v_{\mathtt{i}})\,.$
Thus, we have
$\mathtt{P}(S)$ $=\mathbb{E}_{\sigma_{\mathtt{V}}\sim\nu_{\mathtt{V}}}\mathbb{E}_{\sigma_{\setminus\mathtt{V}}\sim\nu_{\setminus\mathtt{V}}}\Bigg{[}\prod_{1\leq\mathtt{i}\leq\mathtt{t}}\prod_{(i,j)\in E(P_{\mathtt{i}})}\mathtt{h}(\sigma_{i},\sigma_{j})\mid\{\sigma_{u}:u\in\mathtt{V}\}\Bigg{]}$
$=\mathbb{E}_{\sigma_{\mathtt{V}}\sim\nu_{\mathtt{V}}}\Bigg{[}\prod_{1\leq\mathtt{i}\leq\mathtt{t}}\prod_{\{u_{\mathtt{i}},v_{\mathtt{i}}\}=\operatorname{EndP}(P_{\mathtt{i}})}\Big{(}a^{|E(P_{\mathtt{i}})|}+b^{|E(P_{\mathtt{i}})|}\omega(u_{\mathtt{i}},v_{\mathtt{i}})\Big{)}\Bigg{]}\,.$
Similarly we can show that (denote $\Lambda=\{\mathtt{i}_{1},\ldots,\mathtt{i}_{s}\}$)
$\mathtt{Q}(S,H)$ $=\mathbb{E}_{\sigma_{\mathtt{V}}\sim\nu_{\mathtt{V}}}\Bigg{[}\prod_{\mathtt{i}\in[\mathtt{t}]\setminus\Lambda}\prod_{\{u_{\mathtt{i}},v_{\mathtt{i}}\}=\operatorname{EndP}(P_{\mathtt{i}})}\Big{(}a^{|E(P_{\mathtt{i}})|}+b^{|E(P_{\mathtt{i}})|}\omega(u_{\mathtt{i}},v_{\mathtt{i}})\Big{)}\prod_{\mathtt{i}\in\Lambda}\prod_{\{u_{\mathtt{i}},v_{\mathtt{i}}\}=\operatorname{EndP}(P_{\mathtt{i}})}\omega(u_{\mathtt{i}},v_{\mathtt{i}})\Bigg{]}$
$\leq k^{s}\prod_{\mathtt{i}\in\Lambda}a^{-|E(P_{\mathtt{i}})|}\mathbb{E}_{\sigma_{\mathtt{V}}\sim\nu_{\mathtt{V}}}\Bigg{[}\prod_{1\leq\mathtt{i}\leq\mathtt{t}}\prod_{\{u_{\mathtt{i}},v_{\mathtt{i}}\}=\operatorname{EndP}(P_{\mathtt{i}})}\Big{(}a^{|E(P_{\mathtt{i}})|}+b^{|E(P_{\mathtt{i}})|}\omega(u_{\mathtt{i}},v_{\mathtt{i}})\Big{)}\Bigg{]}$
$\leq k^{s}a^{-(|E(S)|-|E(H)|)}\mathtt{P}(S)\,.$
Thus, we have that
$\frac{(\epsilon^{2}\lambda)^{|E(S)|-|E(H)|}\mathtt{Q}(S,H)}{\mathtt{P}(S)}$ $\leq(2k)^{s}(\epsilon^{2}\lambda)^{|E(S)|-|E(H)|}a^{|E(S)|-|E(H)|}$
(4.2)${\leq}(2k)^{3r}(1-\delta/2)^{|E(S)|-|E(H)|}\,.$
Thus, we have (combining the induction hypothesis on $H$)
$|\Xi(S)|$ $\leq(1-\delta/2)^{|E(S)|}n^{-|E(S)|/2}\sum_{1\leq r\leq m}(2k)^{3r}(10(m-r))!\#\{H\subset S:\mathcal{L}(H)=\emptyset\}$
$\leq(1-\delta/2)^{|E(S)|}n^{-|E(S)|/2}\sum_{1\leq r\leq m}(2k)^{3r}(10(m-r))!D^{3r}$
$\leq(10m)!(2kD)^{10m}(1-\delta/2)^{\frac{|E(S)|}{2}}n^{-\frac{|E(S)|}{2}}\,,$
where in the second inequality we use Lemma A.5 with $m_{j}=0$, leading to Item (3). ∎
Now we prove Lemma 4.4 based on Lemma B.5.
Proof of Lemma 4.4.
It is easy to check that $u$ satisfies (4.20). We need to show that
$\sum_{\sigma\in[k]^{n}}\sum_{H\Subset\mathcal{K}_{n},|E(H)|\leq D}u_{\sigma,H}^{2}=\sum_{\begin{subarray}{c}H\Subset\mathcal{K}_{n},\mathcal{L}(H)=\emptyset\\ |E(H)|\leq D\end{subarray}}\Xi(H)^{2}$ (B.44)
is bounded by $O_{\delta}(1)$. Using Lemma B.5, we see that (B.44) is bounded by
$\sum_{\begin{subarray}{c}m_{3},\ldots,m_{D}\geq 0,t\geq 0\\ v\geq 3m_{3}+\ldots+Dm_{D}\end{subarray}}2^{10k+m_{3}+\ldots+m_{D}}k^{10t}n^{-(v+t)}(1-\delta/2)^{v+t}\cdot\mathsf{ENUM}^{\prime}(m_{3},\ldots,m_{D};t)\,,$ (B.45)
where
${}\mathsf{ENUM}^{\prime}(m_{3},\ldots,m_{D};t)=\#\Big{\{}S\Subset\mathcal{K}_{n}:\mathcal{L}(S)=\emptyset,\mathfrak{C}_{k}(S)=m_{k}\mbox{ for }k\leq D,\tau(S)=t\Big{\}}\,.$ (B.46)
Using Lemma A.4 with $H=\emptyset$, we see that
${}\mathsf{ENUM}^{\prime}(m_{3},\ldots,m_{D};t)\leq n^{v}D^{10t}\prod_{3\leq i\leq D}\frac{1}{m_{i}!}\,.$ (B.47)
Plugging (B.47) into (B.45), we get that
[…]
$\leq\sum_{\begin{subarray}{c}m_{3},\ldots,m_{D}\geq 0,t\geq 0\\ v\geq 3m_{3}+\ldots+Dm_{D}\end{subarray}}2^{m_{3}+\ldots+m_{D}}(2^{10}D^{10}k/n)^{10t}(1-\delta/2)^{v}\prod_{3\leq i\leq D}\frac{1}{m_{i}!}$
$\leq[1+o(1)]\sum_{\begin{subarray}{c}m_{3},\ldots,m_{D}\geq 0\\ v\geq 3m_{3}+\ldots+Dm_{D}\end{subarray}}2^{m_{3}+\ldots+m_{D}}(1-\delta/2)^{v}\prod_{3\leq i\leq D}\frac{1}{i^{m_{i}}m_{i}!}$
$\leq O_{\delta,k}(1)\cdot\sum_{m_{3},\ldots,m_{D}\geq 0}2^{m_{3}+\ldots+m_{D}}\prod_{3\leq i\leq D}\frac{1}{i^{m_{i}}m_{i}!}=O_{\delta,k}(1)\,.$
This concludes the desired result. ∎
Appendix C Proof of Proposition 3.3 assuming the low-degree conjecture
Shortly after the submission of an earlier version of this draft, we realized that actually we can prove Proposition 3.3 (and thus proving Theorem 3.2) assuming only the original low-degree conjecture. However, such a proof seems to be more problem dependent so we postpone it here. To be more precise, in this section we will prove Proposition 3.3 assuming the following (truncated) low-degree conjecture.
Conjecture C.1 (Truncated low-degree conjecture).
For “natural” high-dimensional hypothesis testing problems between $\mathbb{P}$ and $\mathbb{Q}$, the following statements hold.
  1. (1)
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=O(1)$ as $n\to\infty$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then there exists a constant $C$ such that no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves strong detection between $\mathbb{P}$ and $\mathbb{Q}$.
  2. (2)
    If $\mathsf{Adv}_{\leq D}\big{(}\tfrac{\mathrm{d}\mathbb{P}^{\prime}}{\mathrm{d}\mathbb{Q}^{\prime}}\big{)}=1+o(1)$ as $n\to\infty$ for some $\mathbb{P}^{\prime},\mathbb{Q}^{\prime}$ such that $\operatorname{TV}(\mathbb{P},\mathbb{P}^{\prime}),\operatorname{TV}(\mathbb{Q},\mathbb{Q}^{\prime})=o(1)$, then there exists a constant $C$ such that no algorithm with running time $n^{D/(\log n)^{C}}$ that achieves weak detection between $\mathbb{P}$ and $\mathbb{Q}$.
Our proof similarly starts by Lemma 3.5. The key observation is that the following weaker version of Lemma 3.6 suffices to deduce the impossibility of partial matching.
Lemma C.2.
Assuming Conjecture C.1, there exists a constant $C$ such that for all $1\leq i\leq n$ and all statistics $\{g_{i,j}=g_{i,j}(A,B):1\leq j\leq n\}$ such that each $g_{i,j}(A,B)$ can be computed in time $n^{D/(\log n)^{C}}$ and
$\sum_{1\leq j\leq n}g_{i,j}(A,B)=o(n)\mbox{ for all }1\leq i\leq n\mbox{ and for all graphs }A,B\mbox{ on }[n]\,,$ (C.1)
it holds that
${}\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\Big{[}\big{(}\mathbf{1}_{\{\pi_{*}(i)=j\}}-g_{i,j}\big{)}^{2}\Big{]}\geq 1-o(1)\,.$ (C.2)
Proof.
Without losing of generality, we may assume that $i=1$ in the following proof. Suppose on the contrary that there are statistics $\{g_{j}=g_{j}(A,B):1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ such that $\sum_{1\leq j\leq n}g_{j}(A,B)=o(n)$ for all graphs $A,B$ and (3.6) holds. Again we may assume that $0\leq g_{i}\leq 1$. Recall (3.7), (3.8) and (3.9). Using Lemma 3.5 and Theorem 2.4, we see that (assuming Conjecture C.1) we have (3.10) still holds. To this end, define $g=g_{1}+\ldots+g_{n}$. Using $0\leq g\leq n$, we see that (3.10) yields that $\mathbb{Q}(g\geq\Omega(n))\geq\Omega(n)$, which violates our assumption (C.1). This implies that (3.6) is impossible and thus completes our proof. ∎
Now we can finish the proof of Item (1) in Proposition 3.3 assuming Conjecture C.1.
Proof of Item (1) in Proposition 3.3 assuming Conjecture C.1.
Suppose on the contrary there are statistics $\{f_{j}:1\leq j\leq n\}$ that can be computed in time $n^{D/(\log n)^{C}}$ satisfying Items (1) and (2) in Definition 3.1 with $\mathbb{E}_{\mathbb{P}_{*}}[f_{\pi_{*}(i)}]\geq 1-c$ for some fixed constant $0.010<c<$ 0.01. We first claim that we may assume that
$f_{i,j}(A,B)\in\{0,1\}\mbox{ and }\sum_{1\leq j\leq n}f_{i,j}(A,B)\in\{0,1\}\mbox{ for all graphs }A,B\mbox{ on }[n]\,.$ (C.3)
In fact, denote
$\mathcal{A}=\Big{(}\cap_{1\leq i,j\leq n}\Big{\{}f_{i,j}(A,B)\in\{0,1\}\Big{\}}\Big{)}\bigcap\Big{(}\cap_{1\leq i\leq n}\Big{\{}\sum_{1\leq j\leq n}f_{i,j}(A,B)=1\Big{\}}\Big{)}\,,$
we may consider the truncated statistics
$f_{i,j}^{\prime}(A,B)=f_{i,j}(A,B)\cdot\mathbf{1}_{\mathcal{A}}\,.$
It is straightforward to check that $\{f_{i,j}^{\prime}(A,B):1\leq i,j\leq n\}$ satisfies (C.3) and we still have
$\mathbb{E}_{\mathbb{P}_{*}}\big{[}f_{i,j}^{\prime}(A,B)\big{]}\geq 1-c\,.$
Thus, we may assume that (C.3) holds without losing of generality. To this end, using (C.3), we see that $f_{j}f_{k}=0$ for all $k\neq j$, and thus
$1=\mathbb{E}_{\mathbb{P}}\big{[}(f_{1}+\ldots+f_{n})^{2}\big{]}=\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}}\big{[}f_{j}^{2}\big{]}\,.$ (C.4)
In addition, we have
$\mathbb{E}_{\mathbb{P}_{*}}\big{[}f_{\pi_{*}(i)}\big{]}=\frac{1}{n}\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\big{[}f_{j}\mid\pi_{*}(i)=j\big{]}\geq c\,.$ (C.5)
Thus, for all $\lambda\in[0,1]$ we have
[…]
Thus, by choosing $\lambda=\lambda(c)$ to be a sufficiently small positive constant we get that
$\sum_{j=1}^{n}\mathbb{E}_{\mathbb{P}_{*}}\Big{[}\big{(}\mathbf{1}_{\{\pi_{*}(i)=j\}}-g_{j}\big{)}^{2}\Big{]}=1-\Omega(1)\mbox{ where }g_{j}=\tfrac{1-\lambda}{n}+\lambda f_{j}\,,$
contradicting to Lemma 3.6. This leads to the desired result. ∎